### v1

# Enzyme-Substrate Complex Formation Modulates Diffusion-Driven Patterning in Metabolic Pathways

Abstract.
We investigate how enzymatic binding kinetics regulate diffusion-driven instabilities in a two-step metabolic pathway. Starting from a mechanistic description in which the substrate reversibly binds to the first enzyme before catalytic conversion, we formulate two reaction–diffusion models: a simplified system with effective kinetics and a non-degenerated system that explicitly includes the transient enzyme–substrate complex. A quasi-steady-state reduction yields a lower-dimensional model that preserves the nonlinear coupling between catalytic turnover and spatial transport while enabling complete linear and weakly nonlinear analyses. We show that explicit complex formation shifts the homogeneous steady state, modifies relaxation dynamics, and substantially alters the size and location of the Turing instability region relative to the simplified model. Numerical simulations agree closely with weakly nonlinear predictions, demonstrating how reversible binding attenuates pattern amplitude and slows the emergence of spatial heterogeneity. These results establish a quantitative link between enzyme–substrate binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as that associated with liquid–liquid phase separation (LLPS). The framework provides a mechanistic route by which biochemical association, dissociation, and catalytic rates jointly regulate the robustness and structure of spatial metabolic patterns, and can be extended to broader classes of compartmentalized biochemical networks.
[…]
F. FARIVAR
[…]
1. Introduction
Understanding the complex interactions in biochemical reaction networks is essential for uncovering the principles of cellular function and metabolic regulation. Among the key features of these systems are reaction-diffusion processes, in which chemical species diffuse through space and interact in nonlinear ways. These interactions can give rise to striking spatial organizations, most notably Turing patterns, which are periodic structures that emerge spontaneously from homogeneous initial conditions when reaction kinetics and diffusion rates satisfy certain instability criteria. Since the pioneering work of Turing [37], reaction-diffusion mechanisms provide a robust mathematical framework for the generation of spatial patterns in biological systems [30, 29]. ranging from chemical oscillators [9], [11] to developmental patterning in multicellular systems [10, 15], .
Metabolic pathways in living organisms provide a particularly rich setting for studying such processes. These pathways consist of sequences of enzymatic reactions that convert substrates into intermediates and final products, often with multiple branches and feedback loops [16], [24], [17]. Spatial aspects of metabolism have received growing attention in recent years, as it has become clear that enzymes and metabolites are not always homogeneously distributed in the cytoplasm. Instead, cells frequently organize enzymes into clusters or into biomolecular condensates through liquid–liquid phase separation (LLPS) [31], [25], [7]. Such mesoscale organization can accelerate metabolic fluxes, buffer fluctuations, and regulate competition at branch points [3], [28], [26], [23, 27]. Liquid‒liquid phase separation is a fundamental mechanism of sepsis [6] From a theoretical perspective, reaction–diffusion models offer a natural framework to investigate how enzyme clustering and LLPS influence the spatial dynamics of metabolism [5, 32], [4, 8, 34, 35, 22]. For instance, liquid-liquid phase separation (LLPS) on cell membranes has been described using reaction-diffusion models, which demonstrate that enzymatic feedback loops may drive out-of-equilibrium dynamics that replicate classical phase separation properties as nucleation, coarsening, and domain creation [35]. A straightforward reaction-diffusion framework called the swarm model has been put out to explain the molecular condensation processes that occur in cells. It demonstrates how these condensates emerge out of equilibrium and may be controlled by nucleation and cooperativity dynamics. It captures important aspects of membraneless organelle development, including nucleation, clustering, diffusion, and particle exchange [8].
In this work, we introduce Models 1, 2, and 3 to investigate the conditions for Turing pattern formation in enzyme-mediated metabolic pathways. The Table 1 summarizing all parameters contained in Models 1 and 2 followed as:
Parameter Description (with units)
[…]
$\alpha_{0}$ Relaxation rate toward $C_{0}^{*}$   [s-1]
[…]
$k_{d_{1}}$ Dissociation rate of $E_{1}S_{0}$   [s-1]
$k_{cat}$ Catalytic turnover rate to $S_{1}$   [s-1]
$\beta$ Degradation rate of $C_{1}$   [s-1]
[…]
Table 1. Summary of parameters used in Models 1 and 2 with physical units.
Unless otherwise stated, all parameters appearing in the models and numerical simulations use the dimensional units listed in Table 1.
The novelty of this work lies in combining mechanistic enzyme–substrate kinetics with diffusion and cross-diffusion to study the emergence of spatial metabolic organization. We introduce a non-degenerated reaction–diffusion formulation that explicitly incorporates reversible enzyme–substrate complex formation and derive a reduced model via a quasi-steady-state approximation that preserves the essential nonlinear feedback between catalysis and spatial transport. By comparing this model with a classical simplified enzymatic pathway, we demonstrate how complex formation shifts the homogeneous steady state, alters relaxation dynamics, and significantly modifies the Turing instability region. To our knowledge, this is the first analysis quantifying how enzymatic binding kinetics reshape diffusion-driven instabilities in two-step metabolic pathways, providing a mechanistic link between enzyme-mediated feedback and mesoscale organization such as liquid–liquid phase separation.
1.1. Model 1: The primary framework.
As a starting point, we consider a simplified two-step metabolic pathway [5]: $S_{0}\xrightarrow{E_{1}}S_{1}\xrightarrow{E_{2}}P,$ where $S_{0}$, $S_{1}$, and $P$ denote substrate, intermediate, and final product, respectively. The substrate $S_{0}$ is converted to the intermediate $S_{1}$ by enzyme $E_{1}$, which in turn is converted to $P$ by enzyme $E_{2}$. Following the phenomenological formulation of Castellana et al. [5], we model the spatio-temporal evolution of the substrate and intermediate as
$\begin{cases}\dfrac{\partial C_{0}}{\partial t}=-\alpha_{0}(C_{0}-C_{0}^{*})-k_{1}n_{1}C_{0}+D_{0}\nabla^{2}C_{0}-d\,C_{0}\nabla^{2}C_{1},\\[6.0pt] \dfrac{\partial C_{1}}{\partial t}=-k_{2}n_{2}C_{1}-\beta C_{1}+k_{1}n_{1}C_{0}+D_{1}\nabla^{2}C_{1}.\end{cases}$ (1.1)
Here $C_{0}$ and $C_{1}$ denote the concentrations of $S_{0}$ and $S_{1}$, while $n_{1}$ and $n_{2}$ represent fixed enzyme levels. The parameter $\alpha_{0}$ drives $C_{0}$ toward its homeostatic value $C_{0}^{*}$, $k_{1}n_{1}$ and $k_{2}n_{2}$ are effective catalytic and consumption rates, and $\beta$ is the degradation rate of $C_{1}$. For $\beta=0$, the intermediate can accumulate indefinitely, eliminating steady-state flux through the pathway.
[…]
This coupling does not depend on the gradient of $C_{1}$ but on its curvature: at locations where $C_{1}$ attains a local maximum, $\nabla^{2}C_{1}<0$ and therefore $d\,C_{0}\nabla^{2}C_{1}>0$, implying that $C_{0}$ increases at such sites. Conversely, $C_{0}$ decreases near local minima ($\nabla^{2}C_{1}>0$). Thus, substrate accumulates preferentially near regions where the intermediate forms concave peaks, producing an effective recruitment mechanism without invoking direct molecular attraction.
Such curvature-dependent transport is consistent with mesoscale biophysical processes. Curvature-sensing proteins in membranes preferentially bind to negatively curved regions [1], and curvature gradients can drive molecular fluxes in cells and tissues [36]. Analogous behavior occurs in LLPS condensates and metabolon-like assemblies, where spatial variations in viscosity, binding affinity, and chemical potential bias metabolite transport [5, 38]. The term $d\,C_{0}\nabla^{2}C_{1}$ therefore provides a minimal and biophysically motivated representation of curvature-driven substrate recruitment.
LLPS proceeds through nucleation, maturation, and functional biochemical activity [19, 2]. Model 1 corresponds to the latter two stages: we assume that a condensate has already formed and maintains approximately constant internal enzyme concentrations $n_{1}$ and $n_{2}$, while the metabolites $C_{0}$ and $C_{1}$ remain mobile and diffuse into, out of, and within the condensed phase. The curvature-driven term $d\,C_{0}\nabla^{2}C_{1}$ captures substrate recruitment into regions of high intermediate accumulation, a hallmark of the mature, functional stage of LLPS-mediated metabolic organization.
Model 1 thus serves as a coarse-grained baseline for studying generic reaction–diffusion feedbacks in enzyme-containing LLPS compartments and provides the foundation upon which mechanistic extensions (Model 2) can be built.
1.2. Model 2: Explicit enzyme–substrate interactions.
Model 1 captures basic spatial coupling but treats catalysis as an effective one-step process and does not resolve the transient enzyme–substrate complex. In cellular environments, however, enzymes bind substrates reversibly before turnover [20], and the mobility of such complexes—especially inside co-localized assemblies such as liquid–liquid phase-separated (LLPS) condensates—may differ substantially from that of free metabolites.
To account for these effects, we introduce an additional dynamical variable representing the enzyme–substrate complex $C_{E_{1}S_{0}}$. The underlying biochemical scheme,
$E_{1}+S_{0}\xrightleftharpoons[K_{d_{1}}]{k_{a_{1}}}E_{1}S_{0}\xrightarrow{k_{cat}}E_{1}+S_{1},$
motivates the extended reaction–diffusion system
$\begin{cases}\dfrac{\partial C_{0}}{\partial t}=-\alpha_{0}(C_{0}-C_{0}^{*})-k_{d_{1}}n_{1}\,C_{E_{1}S_{0}}+\gamma_{0}\nabla^{2}C_{0}-\gamma\,C_{0}\nabla^{2}C_{1},\\[6.0pt] \dfrac{\partial C_{E_{1}S_{0}}}{\partial t}=k_{a_{1}}n_{1}\,C_{0}-(k_{d_{1}}+k_{cat})C_{E_{1}S_{0}}+\gamma_{\mathrm{comp}}\nabla^{2}C_{E_{1}S_{0}},\\[6.0pt] \dfrac{\partial C_{1}}{\partial t}=-(k_{2}n_{2}+\beta)\,C_{1}+k_{cat}\,C_{E_{1}S_{0}}+\gamma_{1}\nabla^{2}C_{1},\end{cases}$ (1.2)
Remark. Equation (1.2) is not intended as a strict microscopic mass-action description. Instead, $C_{E_{1}S_{0}}$ acts as an effective complex-density field summarizing local co-localization of enzyme and substrate—for example within LLPS condensates or metabolon clusters. The terms $k_{a_{1}}n_{1}C_{0}$ and $k_{d_{1}}n_{1}C_{E_{1}S_{0}}$ therefore represent coarse-grained recruitment and release processes characteristic of mesoscale organization [5, 38].
Model 2 extends the phenomenological dynamics of Model 1 by introducing an explicit intermediate while preserving the same basic mechanism: feedback between catalytic activity and spatial metabolite organization. Model 1 provides a minimal two-species description, whereas Model 2 offers a more mechanistic formulation linking enzyme binding, catalytic turnover, and condensate-mediated localization.
1.3. Model 3: Reduced non-degenerate formulation.
While the explicit inclusion of enzyme–substrate complexes enriches the model biologically, it introduces mathematical complications. In particular, the extended three-variable system is degenerate with respect to weakly nonlinear (WNL) analysis, preventing the direct derivation of amplitude equations near the instability threshold. This difficulty arises from the presence of fast binding-unbinding dynamics that render the model stiff and obstruct the perturbative expansion required in WNL theory. To overcome this challenge, we apply a quasi-steady-state reduction by assuming that the enzyme-substrate complex equilibrates rapidly relative to the other variables. This reduction yields a non-degenerate two-variable model (in section 3.2) in which the influence of the complex is encoded in effective reaction terms. Importantly, the reduced model preserves the essential nonlinear feedbacks and LLPS-related cross-diffusion couplings of the modified model, while restoring analytical tractability.
Throughout this work we impose homogeneous Neumann (zero–flux) boundary conditions, for all models and all parameters are primary considered non-negative. These conditions ensure mass conservation and reflect the fact that metabolites and enzyme–substrate complexes do not cross the boundary of the system.
Although the numerical simulations presented in Section 6 are performed in one spatial dimension, the reaction–diffusion mechanism is not restricted to 1D. The same chemical interactions and diffusion laws apply in two and three dimensions, and the Turing conditions are dimension-independent: what changes is only the geometry of the domain and the admissible wave numbers. In higher dimensions, additional pattern morphologies (spots, stripes, labyrinths, hexagons)([33], [13, 14] may occur, but the biochemical interpretation—enzymatic feedback coupled with substrate/intermediate mobility—remains unchanged. Thus, the 1D setting captures the essential instability mechanism while allowing clearer analytical and numerical comparison between the simplified and non-degenerated models.
The remainder of this paper is organized as follows. In Section 2, we analyze the local dynamics of the primary model and derive the conditions for Turing instability driven by diffusion and cross-diffusion effects. In Section 3, we examine the equilibrium structure of the extended three-variable system, showing that it is degenerate. To address this, we employ a quasi–steady-state approximation [20, 33] to obtain a reduced non-degenerate model, for which we investigate both local stability and the onset of Turing instability. Section 4 explores how variations in reaction and diffusion parameters affect the extent of Turing regions in Models 1 and 3. In Section 5, we apply weakly nonlinear (WNL) analysis to the primary and reduced models to derive amplitude equations describing the nonlinear evolution of spatial patterns. These analytical predictions are then compared with numerical simulations in Section 6, confirming the validity of the theoretical results. Finally, Section 7 presents concluding remarks and discusses the broader implications of our findings for understanding spatial regulation in enzyme-mediated and phase-separated biochemical systems.
2. Simplified Two-Step Enzyme Model
In this section, we present linear analysis of the primary model (1.1) around the unique steady state of the system. For all positive non-negative parameters, this system has an unique coexistence steady state $E^{*}=(C_{0}^{E},C_{1}^{E})$ that is:
$C_{0}^{E}=\dfrac{\alpha_{0}C_{0}^{*}}{\alpha_{0}+k_{1}n_{1}},\;\;C_{1}^{E}=\dfrac{k_{1}n_{1}\alpha_{0}*C_{0}^{*}}{(k_{2}n_{2}+\beta)(\alpha_{0}+k_{1}n_{1})},$ (2.1)
Linear analysis of the kinetic around $E^{*}$ demonstrates that $E^{*}$ is always stable since
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-k_{1}n_{1}&&0\\ k_{1}n_{1}&&-\beta-k_{2}n_{2}\end{pmatrix}$, (2.2)
So the characteristic function is obtained for $|\lambda I-J(E^{*})|=\lambda^{2}+\lambda tr(J)+det(J)=0,$ that has two negative eigenvalues $\lambda_{1}=-\alpha_{0}-k_{1}n_{1},\;\;\lambda_{2}=-\beta-k_{2}n_{2}.$
Proposition 2.1.
The system (1.1) does not go under Hopf bifurcation since all parameters are non negative and $tr(J)>0$.
2.1. Turing Analysis
According to the Turing’s statement, a steady state can be unstable in presence of diffusion. In order to investigate Turing instabilities, one needs to linearize whole system at $E^{*}$, which provides:
$\mathcal{D}(E^{*})=\begin{pmatrix}D_{0}&&-dC_{0}^{E}\\ 0&&D_{1}\end{pmatrix}$ (2.3)
and for $W=[C_{0}-C_{0}^{E},C_{1}-C_{1}^{E}]$, we have
$\dot{W}=J(E^{*})W+\mathcal{D}(E^{*})\nabla^{2}W,$
In order to find the solution of the linearized model with replace the system with $W=e^{ikx+\lambda t}$, in which $k$ and $\lambda$ imply wave number and growth rate correspondingly. Therefore, dispersion relation of the model is determined as $|\lambda I-J+k^{2}D|=0$. Thus, the dispersion relation is given by:
$\lambda^{2}+A(k^{2})\lambda+B(k^{2})=0,$ (2.4)
$A(k^{2})=k^{2}(D_{1}+D_{0})+\alpha_{0}+k_{1}n_{1}+\beta+k_{2}n_{2},$ (2.5)
According to the Turing analysis, $E^{*}$ is unstable as one eigenvalue of the dispersion relation is negative, that requires $B(k^{2})<0$ for a range of wave numbers $k_{a},k_{b}]$, where
$B(k^{2})$ $=k^{4}D_{0}D_{1}+k^{2}[-D_{0}(-\beta-k_{2}n_{2})-D_{1}(-\alpha_{0}-k_{1}n_{1})-k_{1}n_{1}dC0_{E}^{*}]$ (2.6)
$+(-k_{1}n_{1}-\alpha_{0})(-\beta-k_{2}n_{2})=b_{1}k^{4}+b_{2}k^{2}+b_{3},$ (2.7)
Turing necessary condition obtains that in the Turing threshold
$B(k_{min}^{2})=0$ (2.8)
which gives
$k^{2}_{min}=\dfrac{-b_{2}}{2b_{1}},$ (2.9)
Since only positive wave numbers are physically meaningful and $b_{1}>0$, therefore, $b_{2}$ must be negative. This provides that critical Turing bifurcation parameter is $d$ and it is determined by:
$d>\dfrac{-D_{0}(-k_{2}n_{2}-\beta)-D_{1}(-k_{1}n_{1}-\alpha_{0})}{k_{1}n_{1}C0_{E}^{*}},$ (2.10)
So to find the critical wave number we replace $d_{c}=\eta_{1}/\eta_{2}+\varepsilon$ where
$\eta_{1}=-D_{0}(-k_{2}n_{2}-\beta)-D_{1}(-k_{1}n_{1}-\alpha_{0}),$ (2.11)
$\eta_{2}=k_{1}n_{1}C0_{E}^{*},$ (2.12)
into (2.8) which obtains that
$\varepsilon=\dfrac{2\sqrt{D_{0}D_{1}(-\alpha_{0}-k_{1}n_{1})(-\beta-k_{2}n_{2})}}{k_{1}n_{1}C0_{E}^{*}},$ (2.13)
and consequently
$k_{min}^{2}=\sqrt{\dfrac{(-\alpha_{0}-k_{1}n_{1})(-\beta-k_{2}n_{2})}{D_{0}D_{1}}}.$ (2.14)
Hence, the system goes to the pattern formation if necessary condition (2.10) is satisfied.
Remark 2.1.
It is easy to check that in absence of cross-diffusion $(d=0)$, they system remain stable since $b_{2}>0$ for any given parameters. So, cross-diffusion term is only key mechanism of emergence of Turing patterns.
3. Modified Model and Quasi–Steady–State Reduction
3.1. Equilibria and Stability Analysis of the Modified Model
To obtain a more realistic representation of the enzyme–substrate interaction within the reaction–diffusion framework, it is necessary to explicitly account for the formation of enzyme–substrate (intermediate) complexes. In enzymatic reactions, the catalytic conversion of the substrate into the product occurs only after the formation of such a complex, which is typically assumed to be irreversible. Therefore, the reaction scheme has been extended by including the additional binding and catalytic steps, leading to the formulation of the modified model (1.2).
We first analyze the equilibria of system (1.2) in the spatially homogeneous case. The system admits two distinct sets of steady states:
  • 1.
    The equilibrium $E_{1}^{*}=(C_{0}^{*},\,0,\,0)$, which exists whenever $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})\neq 0$.
  • 2.
    Supposing $\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}C_{E_{1}S_{0}}^{*}>0$, the family of equilibria
    $E_{2}^{*}=\Big(\dfrac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}C_{E_{1}S_{0}}^{*}}{\alpha_{0}},\,C_{E_{1}S_{0}}^{*},\,\dfrac{k_{cat}C_{E_{1}S_{0}}^{*}}{k_{2}n_{2}+\beta_{2}}\Big),$
    which appears in the degenerate case $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ and corresponds to a continuous line of equilibria.
Indeed, the degeneracy arises because the binding/unbinding dynamics are fast relative to metabolite dynamics, giving one zero eigenvalue in the Jacobian and producing a line of equilibria. This prevents standard WNL expansion, motivating the QSSA.
Theorem 3.1.
Let $E_{1}^{*}$ denote the equilibrium point of system (1.2). Then:
  1. (1)
    $E_{1}^{*}$ is locally asymptotically stable if $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})<0$.
  2. (2)
    $E_{1}^{*}$ is unstable if $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})>0$.
Proof.
Linearizing system (1.2) around the equilibrium $E_{1}^{*}$ yields the Jacobian matrix
$J(E_{1}^{*})=\begin{bmatrix}-\alpha_{0}&-k_{d_{1}}n_{1}&0\\[5.69054pt] 0&k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})&0\\[5.69054pt] 0&k_{cat}&-(k_{2}n_{2}+\beta_{2})\end{bmatrix}$. (3.1)
The eigenvalues of $J(E_{1}^{*})$ are $\lambda_{1}=-\alpha_{0}$, $\lambda_{2}=-(k_{2}n_{2}+\beta_{2})$, and $\lambda_{3}=k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})$. Since $\alpha_{0}>0$ and $(k_{2}n_{2}+\beta_{2})>0$, the first two eigenvalues are negative. The sign of $\lambda_{3}$ therefore determines the local stability of $E_{1}^{*}$. If $\lambda_{3}<0$, all eigenvalues are negative and the equilibrium is locally asymptotically stable. Conversely, if $\lambda_{3}>0$, one eigenvalue is positive and $E_{1}^{*}$ becomes unstable. This completes the proof. ∎
The above theorem shows that the stability of the equilibrium $E_{1}^{*}$ in the modified (non-degenerated) enzyme model depends solely on the relative magnitudes of the association, dissociation, and catalytic rates. When the binding rate $k_{a_{1}}n_{1}$ exceeds the sum of the dissociation and catalytic rates $(k_{d_{1}}+k_{cat})$, the feedback becomes self-amplifying, leading to an unstable steady state. Conversely, when the effective dissociation and catalytic losses dominate, the equilibrium remains locally asymptotically stable.
Remark 3.2.
At the critical value $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$, the Jacobian matrix possesses two negative eigenvalues and one zero eigenvalue, and the system exhibits a one-dimensional manifold of equilibria. Each point on this line is non-hyperbolic and neutrally stable along the direction of the equilibrium manifold.
Theorem 3.3.
Degenerate case: line of equilibria and neutral direction] Assume $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ in the modified model (1.2). Then:
1. The set $\mathcal{M}\;=\;\Big\{\,E_{2}^{*}(\xi)\;=\;\Big(\frac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}\xi}{\alpha_{0}}\;,\;\xi\;,\;\frac{k_{cat}}{k_{2}n_{2}+\beta_{2}}\,\xi\Big)\;:\;\xi\in\mathcal{I}\Big\}$ (with $\mathcal{I}$ chosen to enforce nonnegativity) is a one–dimensional manifold of equilibria.
2. For every $E\in\mathcal{M}$, the Jacobian $J(E)$ has eigenvalues $\lambda_{1}=-\alpha_{0}<0,\;\lambda_{2}=-(k_{2}n_{2}+\beta_{2})<0,\;\lambda_{3}=0.$ In particular, each $E\in\mathcal{M}$ is non-hyperbolic, with a neutral direction tangent to $\mathcal{M}$ and exponential stability in the two transverse directions.
Proof.
Setting the time derivatives in system (1.2) to zero and imposing the degeneracy condition $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$ removes one scalar constraint among the steady–state equations. Solving the resulting algebraic system yields the one–parameter family of equilibria $E_{2}^{*}(\xi)$ described in (1); hence every point on $\mathcal{M}$ is an equilibrium.
Evaluating the Jacobian matrix (3.1) at the degenerate condition $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ and at $E_{2}^{*}(\xi)$ gives three eigenvalues: $\lambda_{1}=-\alpha_{0}<0$, $\lambda_{2}=-(k_{2}n_{2}+\beta)<0$, and $\lambda_{3}=0$. A tangent vector to $\mathcal{M}$ at $E_{2}^{*}(\xi)$ is $v=\frac{d}{d\xi}E_{2}^{*}(\xi)=\Big(-\tfrac{k_{d_{1}}n_{1}}{\alpha_{0}},\,1,\,\tfrac{k_{cat}}{k_{2}n_{2}+\beta}\Big)^{\!\top}.$ Direct substitution shows that $J(E_{2}^{*}(\xi))\,v=0$, so $v$ spans the nullspace of the Jacobian, confirming that the zero eigenvalue corresponds to the direction tangent to the equilibrium manifold $\mathcal{M}$.
According to the Center Manifold Theorem ([21]), a zero eigenvalue of the Jacobian gives rise to a one-dimensional local center manifold $W^{c}_{\mathrm{loc}}(E)$ tangent to the associated eigenvector. In the present case, this manifold coincides locally with the equilibrium line $\mathcal{M}$. The remaining two eigenvalues are strictly negative, and by the Stable Manifold Theorem, they generate a two–dimensional local stable manifold $W^{s}_{\mathrm{loc}}(E)$ on which perturbations decay exponentially to $\mathcal{M}$. Consequently, each $E\in\mathcal{M}$ is a non–hyperbolic equilibrium, characterized by exponential contraction in the directions transverse to $\mathcal{M}$ and neutral (center) stability along the manifold itself. ∎
3.2. Equilibrium Analysis and Derivation of the Non-degenerated Model
To obtain the equilibrium concentrations $C_{0}^{*}$, $C_{1}^{*}$, and $C_{E_{1}S_{0}}^{*}$, we first note that from the first equation of system (1.2) one can express $C_{0}=C_{0}^{*}-\dfrac{k_{d_{1}}n_{1}\,C_{E_{1}S_{0}}}{\alpha_{0}}.$ From the second equation, under the simplifying assumption $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$, the model becomes degenerate. Solving the third equation for $C_{1}$ gives $C_{1}=\dfrac{k_{cat}}{k_{2}n_{2}+\beta}\,C_{E_{1}S_{0}}.$
Since the system is thus degenerate, to obtain a reduced model with independent variables we employ the quasi–steady–state approximation (QSSA). This assumption states that the enzyme–substrate complex $C_{E_{1}S_{0}}$ relaxes much faster than the concentrations of $C_{0}$ and $C_{1}$, i.e. $\dfrac{\partial C_{E_{1}S_{0}}}{\partial t}\approx 0.$ Hence, from the second equation of system (1.2) we can write $C_{E_{1}S_{0}}=\dfrac{k_{a_{1}}n_{1}C_{0}+D\nabla^{2}C_{E_{1}S_{0}}}{k_{d_{1}}+k_{cat}}.$
Assuming that the diffusive contribution $D\nabla^{2}C_{E_{1}S_{0}}$ is negligible compared with the reaction term $k_{a_{1}}n_{1}C_{0}$, we obtain the approximation
$C_{E_{1}S_{0}}\approx\dfrac{k_{a_{1}}n_{1}}{k_{d_{1}}+k_{cat}}\,C_{0}.$ (3.2)
Substituting (3.2) into the remaining equations of system (1.2) yields the following reduced or non-degenerated model:
$\begin{cases}\dfrac{\partial u_{0}}{\partial t}=-\alpha_{0}\big(u_{0}-u_{in}^{*}\big)-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}}\,u_{0}+\gamma_{0}\nabla^{2}u_{0}-\gamma\,u_{0}\,\nabla^{2}u_{1},\\[5.69054pt] \dfrac{\partial u_{1}}{\partial t}=-(k_{2}n_{2}+\beta)\,u_{1}+\dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}\,u_{0}+\gamma_{1}\nabla^{2}u_{1}.\end{cases}$ (3.3)
The non-degenerated system (3.3) retains the same structure as the primary model (1.1), but with modified effective coefficients that encapsulate the influence of the enzyme–substrate complex through the QSSA. The variable $C_{E_{1}S_{0}}$ has been successfully eliminated, and the resulting model now involves only the substrate and intermediate/product concentrations $(C_{0},C_{1})$. This formulation allows for direct comparison with the simplified model (1.1), while maintaining the mechanistic effects of enzyme–substrate binding and catalytic conversion.
3.3. Linear Analysis of enzyme-substrate complex
The model (3.3) has an unique coexistence steady state as:
$u_{0}^{*}=\dfrac{\alpha_{0}u_{in}^{*}(k_{d_{1}}+k_{cat})}{\alpha_{0}(k_{d_{1}}+k_{cat})+k_{a_{1}}k_{d_{1}}n_{1}^{2}},\;\;u_{1}^{*}=\dfrac{k_{a_{1}}k_{cat}n_{1}\alpha_{0}u_{in}^{*}}{(k_{d_{1}}+k_{cat})(k_{a_{1}}k_{d_{1}}n_{1}^{2})}$ (3.4)
This model contains all positive parameters. Linearizing the model at $E^{*}=(u_{0}^{*},u_{1}^{*})$ gives
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n^{2}}{k_{d_{1}}+k_{cat}}&&0\\ \dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}&&-\beta-k_{2}n_{2}\end{pmatrix},\;\;\;\mathcal{D}(E^{*})=\begin{pmatrix}\gamma_{0}&&-\gamma u_{0}^{*}\\ 0&&\gamma_{1}\end{pmatrix}$ (3.5)
This system is locally stable at $E^{*}$ simultaneously. And it admits Turing instability in presence of diffusion with necessary condition
$\gamma>\dfrac{-\gamma_{0}(-k_{2}n_{2}-\beta)-\gamma_{1}(-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}})}{\dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}u_{0}^{*}},$ (3.6)
So $\gamma$ is Turing bifurcation parameter. Moreover, critical wave number is respectively determined
$k_{min}^{2}=\sqrt{\dfrac{(\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}})(-k_{2}n_{2}-\beta)}{\gamma_{0}\gamma_{1}}},$
Figure 1. Phase-plane comparison of the simplified and non-degenerated enzyme-pathway models. Arrows represent vector fields (dark gray: primary model; light gray: non-degenerated model). Red/blue and magenta/green curves denote the corresponding nullclines, whose intersections yield homogeneous steady states (dots). Trajectories from identical initial conditions are shown in magenta and green. The non-degenerated formulation shifts the steady state toward higher intermediate concentration and exhibits slower relaxation, consistent with nonlinear feedback between catalysis and compartmentalization.
We have investigated numerically necessary conditions of appearance of Turing instabilities of both models (1.1) and (3.3) of a given set of parameters (see (Figure 2)). In this figure, plot of $B(k^{2})$ demonstrate as bifurcation parameters $d$ of (1.1) and $\gamma$ of (3.3) crosses their critical values, necessary conditions (2.10) and (3.6) are satisfied and correspondingly for both models $B(k^{2})$ admit negative value and respectively $\lambda(k^{2})$ admit positive values, which states appearance of pattern formation.
Figure 2. Plots of polynomials $B(k^{2})$ of two models (3.3) (red plots) and (1.1) (black plots) (right figure). Plots of dispersion relation polynomials $\lambda^{2}(k^{2})$ (left figure). Given parameters are $\alpha_{0}=0.01(s^{-1})$, $k_{1}=0.05(\mu M^{-1}s^{-1})$, $k_{2}=0.02(\mu M^{-1}s^{-1})$, $k_{d_{1}}=0.01s^{-1}$, $k_{a_{1}}=0.03(\mu M^{-1}s^{-1})$, $k_{cat}=0.1s^{-1}$, $\beta=0.01(s^{-1})$, $C_{0}^{*}=0.1(\mu Ms^{-1})$, $n_{1}=10(\mu M)$, $n_{2}=5(\mu M)$, $\gamma_{0}=0.1(\mu m^{2}s^{-1})$, $\gamma_{1}=0.01(\mu m^{2}s^{-1})$. where in this figures $d_{c}$ and $\gamma_{c}$ are Turing bifurcation parameter of model (1.1) and (3.3) respectively.
A qualitative comparison between the kinetic dynamics of the simplified model (1.1) and the reduced non-degenerated model (3.3) is shown in Figure 1. The phase–plane analysis highlights several key differences: (i) the nullclines of the non-degenerated model intersect at a steady state with higher intermediate concentration, (ii) trajectories relax more slowly toward equilibrium, reflecting the additional nonlinear feedback introduced by enzyme–substrate binding, and (iii) the vector fields of the two systems display distinct curvature, indicating that the mechanistic correction introduced by the QSSA reduction significantly alters the underlying kinetics.
4. Turing region
To understand how parameters influence on appearance of Turing instability of both models, we need to obtain Turing regions. In order to achieve this, we need to consider positivity conditions of both $E^{*}$, stability conditions in absence of diffusion. According to the (2.1) and (3.4) coexistence steady states are always positive for choosing any positive values of parameters. Moreover, it is locally stable for choosing any positive values of parameters. Thus, there are only Turing conditions (2.10), and (3.6). According to these conditions, we investigate variation of parameters in plane $(n_{1},d),\;(n_{1},\gamma)$. Figure 2, demonstrates that for the given data set, the bifurcation parameters of models (1.1) and (3.3) are $d_{c}=2.7190$ and $\gamma_{c}=5.4035$ correspondingly.
4.0.1. Influence of Intermediate Diffusion on the Stability of Enzyme–Substrate Spatial Organization in model (1.1)
To investigate the influence of diffusion on pattern formation and aggregation, we examined how variations in the diffusion coefficients of both the substrate ($C_{0}$) and the intermediate ($C_{1}$) affect the emergence of spatial structures. Our analysis reveals that increasing either $D_{0}$ or $D_{1}$ leads to a progressive reduction of the Turing instability region. As diffusion becomes more efficient, the concentration gradients of both chemical species are rapidly smoothed out, preventing the formation of localized substrate-enriched zones. Consequently, the spatial structures (lobes or peaks) that would otherwise arise from enzyme–substrate feedback are flattened, and the system tends toward a spatially homogeneous steady state. This stabilizing effect of molecular mobility is clearly visible in Figure 3, where the Turing domain diminishes with increasing $D_{0}$ and $D_{1}$.
Figure 3. Variation of Turing regions in plane $(n_{1},d)$ as the parameters $D_{0}$ (left plot) and $D_{1}$ (right plot) increase. Red dots in the plots indicate $d_{c}$ threshold for each given $D_{0}$ or $D_{1}$. Parameters have chosen like Figure 2.
4.0.2. Influence of Intermediate Diffusion on the Stability of Enzyme–Substrate Spatial Organization in model (3.3)
The analysis of Model (3.3) reveals that the diffusion of the intermediate plays a crucial role in regulating the onset of spatial organization. In Figure 4, as the diffusivity of the intermediate increases, the Turing region progressively shrinks and the instability threshold $\gamma_{c}(n_{1})$ shifts toward higher values of the cross-diffusion parameter. This behavior indicates that enhanced intermediate mobility tends to equalize its spatial distribution, thereby weakening the gradients necessary to sustain localized substrate–intermediate interactions. From a biochemical perspective, this corresponds to a reduction in spatial compartmentalization: the mobile intermediate rapidly redistributes throughout the domain, preventing the formation of enzyme–substrate-enriched zones and driving the system toward a spatially homogeneous steady state.
In both the primary and non-degenerated models, increasing the substrate diffusivity $D_{i},\gamma_{i}$ reduces the Turing region, indicating that higher substrate mobility stabilizes the homogeneous state. However, in the non-degenerated model, where the cross-diffusion coupling depends explicitly on substrate concentration and reaction kinetics, this stabilizing influence is stronger. Thus, while both models exhibit diffusion-mediated suppression of pattern formation, the mechanistic coupling in the non-degenerated model amplifies the homogenizing effect.
Figure 4. Variation of Turing regions in plane $(n_{1},\gamma)$ as the parameters $\gamma_{0}$ (left plot) and $\gamma_{1}$ (right plot) increase. Red dots in the plots indicate $\gamma_{c}$ threshold for each given $\gamma_{0}$ or $\gamma_{1}$. Parameters have chosen like Figure 2.
4.0.3. Influence of Kinetic Parameters on the Stability of Enzyme–Substrate Spatial Organization in Model (3.3)
Figures 5 illustrate the influence of the kinetic parameters $k_{a_{1}}$, $k_{d_{1}}$, and $k_{cat}$ on the onset of diffusion-driven instability. In all cases, the system displays the characteristic U-shaped Turing boundary, but the extent and position of the unstable region are strongly affected by the reaction rates. Increasing the association rate $k_{a_{1}}$ enlarges the Turing region and shifts it toward higher $n_{1}$, indicating that stronger substrate–enzyme binding enhances the nonlinear coupling responsible for spatial self-organization. Conversely, increasing the dissociation rate $k_{d_{1}}$ diminishes the Turing domain, showing that faster complex dissociation stabilizes the homogeneous steady state. A similar expansion of the Turing region is observed for higher catalytic rates $k_{cat}$, where a more efficient turnover of the intermediate amplifies local concentration gradients and reinforces the enzyme–substrate feedback. Overall, strong association and high catalytic activity favor the emergence of spatial structures, whereas rapid dissociation weakens the coupling and suppresses pattern formation.
Figure 5. Variation of Turing regions in plane $(n_{1},\gamma)$ as the parameter $k_{cat},\;k_{d_{1}},\;k_{a_{1}}$ increases res. Parameters have chosen like Figure 2.
Figure 6. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left plot) as the parameter $\beta$ increases. Parameters have chosen like Figure 2.
Figure 7. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left) as the parameter $\alpha$ increases. Parameters have chosen like Figure 2.
The figures 6 and 7 investigate how Turing instability regions shift under variations of key parameters, namely $\beta$ and $\alpha_{0}$, in the $(n_{1},d)$ and $(n_{1},\gamma)$ planes. In both the $(n_{1},d)$ and $(n_{1},\gamma)$ plots, increasing $\beta$ leads to higher critical values for $d$ and $\gamma$, respectively. This indicates that larger degradation or reaction rates of the inhibitor suppress pattern formation, requiring stronger diffusion or cross-diffusion to induce Turing instability. Conversely, the $(n_{1},d)$ and $(n_{1},\gamma)$ plots with varying $\alpha_{0}$ show that decreasing $\alpha_{0}$ significantly broadens the Turing region. In $(n_{1},d)$ plane, when $\alpha_{0}\rightarrow 0$, the threshold for $d$ approaches zero, suggesting that slower decay of the activator greatly facilitates the onset of instability.
In summary, the figures illustrate that the Turing space is highly sensitive to kinetic parameters. Smaller values of $\beta$ and $\alpha_{0}$ enlarge the admissible region in the parameter space where spatial patterns can form. This implies that systems with slower degradation or decay kinetics are more susceptible to Turing-type pattern formation. These insights are particularly relevant for enzymatic and reaction-diffusion models in biological systems, where parameter tuning can drive or suppress morphogenetic processes.
Figure 8. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left plot). Variation of Turing regions in plane $(n_{2},d)$ and $(n_{2},\gamma)$ (right plot). Parameters have chosen like Figure 2.
The plotted critical curves Figure 8 highlight how the onset of pattern formation depends nonlinearly on enzyme levels, with broader instability regions for low-to-moderate $n_{1}$, and increasingly suppressed Turing regions with high $n_{2}$. We analyze the critical diffusion parameters $d_{c}$ and $\gamma_{c}$ associated with Turing instability in two enzyme-based reaction-diffusion models. In the $n_{1}$-parameter space, both curves exhibit a non-monotonic, U-shaped behavior. This indicates that intermediate values of $n_{1}$ (the enzyme responsible for the first reaction step) are most conducive to pattern formation, requiring minimal diffusion or cross-diffusion to trigger instability. At both low and high $n_{1}$, the system becomes more stable, and higher values of $d_{c}$ and $\gamma_{c}$ are needed to induce patterns.
In contrast, the plots against $n_{2}$ show that both $d_{c}$ and $\gamma_{c}$ increase monotonically with enzyme concentration. This implies that the second enzyme, involved in downstream processes, stabilizes the system by diminishing feedback necessary for pattern formation. The widening gap between the two curves also suggests that the effect of cross-diffusion becomes increasingly significant as $n_{2}$ increases, reinforcing the role of downstream reactions in regulating spatial instabilities.
5. Weakly nonlinear Analysis of models
In this section we perform a weakly nonlinear analysis for the model (1.1) to predict the amplitude of the pattern near the Turing threshold. The weakly nonlinear analysis is based on the multiple scale methods [12]. The similar strategy has applied on the model (3.3). Since near to the bifurcation the amplitude of the pattern (diffusion-driven instability) has slow temporal scale, then a new temporal scale is defined.
The solution of the original system is written as a weakly nonlinear expansion in the small control parameter $\varepsilon$. We choose $\varepsilon^{2}=\dfrac{d-d_{c}}{d_{c}}.$
The slow scale is obtained from the linear analysis: it is easy to prove that $\lambda\sim\varepsilon^{2}$ and, since the growth rate of the perturbation is proportional to the $exp(\lambda t)$, the slow time scale $T$ is order $\varepsilon^{2}$. Therefore, close to the threshold we separate the fast time $t$ and slow time $T=\varepsilon^{2}t$, so that time derivative is obtained as $\partial_{t}\rightarrow\partial_{t}+\varepsilon^{2}\partial_{T}$.
We separate the linear part from the nonlinear part:
$\partial_{t}\bm{w}=\mathcal{L}^{d_{c}}\bm{w}+\mathcal{Q}_{D}^{d_{c}}(\bm{w},\bm{\nabla^{2}w}),\qquad\bm{w}=[C_{0}-C_{0}^{E},C_{1}-C_{1}^{E}],$ (5.1)
and linear operator is defined as $\mathcal{L}^{d_{c}}=\Gamma J+\mathcal{D}^{d_{c}}\nabla^{2},$ where $\mathcal{D}$ and J{J} have defined in (2.2 and 2.3) and nonlinear operators $\mathcal{Q}_{D}^{d_{c}}(\bm{x},\bm{y})$ are introduced as: $\bm{x}=(x_{C_{0}},x_{C_{1}}),\;\bm{y}=(y_{C_{0}},y_{C_{1}})$,
$\mathcal{Q}_{D}(\bm{x},\bm{y})=\begin{bmatrix}-dx_{c_{0}}y_{c_{1}}\\ 0\end{bmatrix}$, (5.2)
And moreover the bifurcation parameter and the solution are expanded asymptotically
$d=d_{c}+\varepsilon^{2}d^{(2)}+O(\varepsilon^{4}),$ (5.3)
$\bm{w}=\varepsilon\bm{w}_{1}+\varepsilon^{2}\bm{w}_{2}+\varepsilon^{3}\bm{w}_{3}+O(\varepsilon^{4}),$ (5.4)
Therefore, the diffusion matrix is given in terms of perturbation parameters as:
$\mathcal{D}=\begin{bmatrix}D_{0}&-d_{c}C_{0}^{E}\\ 0&D_{1}\end{bmatrix}+\varepsilon^{2}\begin{bmatrix}0&-d^{(2)}C_{0}^{E}\\ 0&0\end{bmatrix}+O(\varepsilon^{4})$ (5.5)
and consequently
$\mathcal{Q}_{D}(\bm{w},\nabla^{2}\bm{w})$ $=\varepsilon^{2}d_{c}\mathcal{Q}_{D}(\bm{w}_{1},\nabla^{2}\bm{w}_{1})+\varepsilon^{3}d_{c}\{\mathcal{Q}_{D}(\bm{w}_{1},\nabla^{2}\bm{w}_{2})$ (5.6)
$+\mathcal{Q}_{D}(\bm{w}_{2},\nabla^{2}\bm{w}_{1})\}+\varepsilon^{4}d^{(2)}\mathcal{Q}_{D}(\bm{w}_{2},\nabla^{2}\bm{w}_{2})+O(\varepsilon^{5}),$ (5.7)
Now we replace all expansion in (5.3), (5.5),(5.4), and (5.6) and sort according to the order of $\varepsilon$. At $O(\varepsilon)$,we obtain the following linear problem:
$\mathcal{L}^{d_{c}}\bm{w}_{1}=0,\quad\bm{w}_{1}=A(T){\rho}\cos(k_{c}x),$ (5.8)
such that satisfying the homogeneous boundary conditions where $\rho$} belongs to the $ker(J-k^{2}_{c}\mathcal{D}^{d_{c}})$. In this stage the $A(T)$, the amplitude of the pattern is arbitrary and the vector $\rho$ is considered constant such whose normalization is
${\rho}=\begin{bmatrix}1\\ \rho_{2}\end{bmatrix}$, (5.9)
where $\rho_{2}=-\dfrac{-\alpha_{0}-k_{1}n_{1}-k_{c}^{2}D_{0}}{dC_{0}^{E}k_{c}^{2}}$.
Moreover, at $O(\varepsilon^{2})$ there is this linear equation which must be solved:
$\mathcal{L}^{d_{c}}\textbf{w}_{2}=\bm{F},$
According to the Fredholm alternative theorem, this equation has a solution if and only if $\langle\bm{F},\psi\rangle=0,$ where $\psi^{*}$ is defined at
$\psi=\begin{bmatrix}1\\ R_{1}^{*}\end{bmatrix}\cos(k_{c}x)=\psi^{*}\cos(k_{c}x),$
and $\langle.,.\rangle$ implied the scalar product in $\bm{L}^{2}(0,\dfrac{2\pi}{k^{c}})$ and $\psi^{*}\in Ker(J-k^{2}_{c}D^{d_{c}})^{\dagger}$ where $\dagger$ shows transpose of complex conjugate of the matrix.
In particular, $\bm{F}=-\dfrac{1}{4}A^{2}\sum_{i=0,2}\mathcal{M}_{i}(\rho,(\rho)\cos(ik_{c}x)$, in which $\mathcal{M}_{i}(\rho,\rho)=-i^{2}k_{c}^{2}\mathcal{Q}_{D}^{d_{c}}(\rho,\rho)$. Hence, the vector $\bm{w}_{2}$ is defined as
$\bm{w}_{2}=A^{2}(\bm{w}_{20}+\bm{w}_{22}\cos(2k_{c}x))$ so that $\mathcal{L}_{i}^{d_{c}}\bm{w}_{2i}=-\dfrac{1}{4}\mathcal{M}_{i}(\rho,\rho),\;i={0,2}$ and $\mathcal{L}_{i}=-i^{2}k^{2}_{c}D^{d_{c}}$.
In following, at $O(\varepsilon^{3})$ we obtain the linear problem
$\mathcal{L}^{d_{c}}\bm{w}_{3}=\bm{G},$ (5.10)
where $\bm{G}=\left(\dfrac{dA}{dT}\rho+A\bm{G}_{1}^{(1)}+A^{3}\bm{G}_{1}^{(3)}\right)\cos(k_{c}x)+A^{3}\bm{G}_{3}\cos(3k_{c}x),$ in which
$\bm{G}_{1}^{(1)}=k^{2}_{c}\begin{bmatrix}0&d^{(2)}C_{0}^{E}\\ 0&0\end{bmatrix}\rho,$
$\bm{G}_{1}^{(3)}=-k_{c}^{2}\mathcal{Q}_{D}(\rho,\bm{w}_{22})-\dfrac{1}{2}k_{c}^{2}\mathcal{Q}_{D}(\bm{w}_{22},\rho),$
$\bm{G}_{3}=3k^{2}_{c}\mathcal{Q}_{D}^{d_{c}}(\rho,\bm{w}_{22}),$
Finally, by applying the solvability condition, we obtain the Stuart-Landau equation for the amplitude
$\dfrac{\partial A}{\partial T}=\sigma A-LA^{3},$
In addition, solvability of the equation (5.10) depends on $\langle\bm{G},\psi\rangle=0$, whose coefficients are given by
$\sigma=-\dfrac{\langle\bm{G}_{1}^{(1)},\psi\rangle}{\langle\rho,\psi\rangle},\;\;\bm{L}=\dfrac{\langle\bm{G}_{1}^{(3)},\psi\rangle}{\langle\rho,\psi\rangle},$
In the Stuart-Landau equation the coefficient $\sigma$ is always positive while L could be either negative or positive, corresponding to a subcritical or supercritical bifurcation.
The nontrivial solution of the amplitude equation is
$\bm{A}_{\infty}=\sqrt{\dfrac{\sigma}{L}},$ (5.11)
which requires $L>0$, and therefore the result of this analysis is valid only for supercritical bifurcations.
Therefore, the asymptotic behavior of the solution is given by weakly nonlinear analysis of $O(\varepsilon^{3})$ is:
$\bm{W}=\varepsilon A\rho cos(k_{c}x)+\varepsilon^{2}A^{2}[\bm{w}_{20}+\bm{w}_{22}cos(2k_{c}x)]+O(\varepsilon^{3}).$ (5.12)
6. Numerical Results
In this section, we exhibit numerical results. Simulations were performed using the dimensional parameters reported in Table 1. The numerical results clearly show that the primary model (Model 1) and the non-degenerated model (Model 3) exhibit qualitatively similar Turing patterns, but with important quantitative differences in both amplitude and relaxation dynamics. In Model 1, the substrate–intermediate feedback is direct: the production of the intermediate $C_{1}$ depends instantaneously on the local availability of substrate $C_{0}$, and the curvature-driven cross-diffusion term $dC_{0}\nabla^{2}C_{1}$ reinforces this coupling by transporting substrate toward regions where $C_{1}$ accumulates. As a result, the effective positive feedback loop between $C_{0}$ and $C_{1}$ is strong, leading to patterns with relatively large amplitude and fast convergence toward the stationary profile (Figure 9).
Figure 9. Turing Pattern for given parameters: $n_{1}=3.5,\;n_{2}=1,\;k_{1}=k_{2}=2,\;\alpha_{0}=1,\;C_{0}^{*}=7,\;\beta=3,\;D_{0}=2,\;D_{1}=2.5$ that provides $C_{0}^{E}=0.87,\;C_{1}^{E}=1.22,\;k_{c}=2.8,\;d_{c}=9.51$.
Figure 10. Turing Pattern for given parameters: $n_{1}=3.5,\;n_{2}=1,\;k_{a_{1}}=5,\;k_{cat}=8,\;k_{d_{1}}=2.5;k_{2}=2,\;\alpha_{0}=1,\;C_{0}^{*}=7,\;\beta=3,\;D_{0}=2,\;D1=2.5$ that provides $u_{0}^{E}=0.44,\;u_{1}^{E}=1.19,\;k_{c}=2.5,\;d_{c}=14.76$.
In contrast, the non-degenerated model incorporates an explicit enzyme–substrate interaction, in which $S_{0}$ must first bind to the enzyme $E_{1}$ to form the complex $E_{1}S_{0}$ before being converted into the intermediate $S_{1}$. This reversible binding acts as a biochemical buffer: part of the substrate is temporarily stored in the complex, and the effective rate at which $S_{0}$ contributes to the production of $S_{1}$ is reduced to $\frac{k_{a_{1}}k_{\mathrm{cat}}n_{1}}{k_{d_{1}}+k_{\mathrm{cat}}}\;<\;k_{1}n_{1}.$ Consequently, the feedback between substrate and intermediate is weaker, and perturbations grow more slowly. This mechanistic buffering is visible in both the phase–plane trajectories and the numerical simulations: patterns in the non-degenerated model have smaller amplitude, smoother spatial profiles, and longer relaxation times (Figure 10).
Furthermore, the modified cross-diffusion structure of Model 3, where the coupling depends explicitly on the substrate concentration $u_{0}$, reduces the sensitivity of the system to spatial gradients. This further stabilizes the homogeneous state and explains why the Turing region of Model 3 is narrower and shifted toward higher diffusion or cross-diffusion values compared with Model 1.
Figure 11. Numerical results vs WNL of model 1 for given parameters: as Figure 9 and $\varepsilon^{2}=0.003$.
Figure 12. Numerical results vs WNL of model 2 for given parameters: as Figure 10 and $\varepsilon^{2}=0.0001$.
The excellent agreement between numerical simulations and the weakly nonlinear predictions (Figures 11-12) confirms that these differences arise from fundamental kinetic mechanisms rather than numerical artifacts. Overall, the numerical results demonstrate that explicit enzyme–substrate binding attenuates the strength of the feedback necessary for Turing instability, leading to weaker and more finely regulated spatial patterns. Biologically, this means that enzymatic complex formation acts as a natural buffering process that prevents excessive accumulation of intermediates and ensures more stable metabolic organization.
7. Conclusion
This work examined how enzymatic reaction kinetics and diffusion-driven transport interact to generate spatial metabolic organization. The linear stability and Turing analyses showed that diffusion and curvature-driven cross-diffusion alone can destabilize otherwise stable homogeneous steady states, producing spatial patterns through local enzyme–substrate feedback. Incorporating explicit complex formation introduces additional kinetic structure that modifies both the reaction balance and the effective diffusion terms, thereby shifting the Turing threshold and altering the size and location of the instability region. The quasi-steady-state reduction yields a non-degenerate model that preserves the essential nonlinear coupling between catalysis and compartmentalization while enabling a complete weakly nonlinear (WNL) analysis.
Theoretical predictions from the WNL framework were strongly supported by numerical simulations, which showed excellent agreement in both pattern amplitude and spatial profile. The comparison between the simplified and non-degenerate models revealed systematic kinetic and spatial differences: explicit enzyme–substrate binding shifts the equilibrium toward higher intermediate concentrations, slows relaxation toward steady state, and produces spatial patterns of smaller amplitude. These results highlight the buffering role of reversible complex formation and demonstrate how enzymatic binding kinetics regulate the strength of the feedback loop that drives Turing instability.
Beyond their mathematical implications, these findings have a natural interpretation in the context of liquid–liquid phase separation (LLPS). Inside biomolecular condensates, enzymes and metabolites experience altered mobility, molecular crowding, and transient binding events that modulate both catalytic flux and spatial organization. The simplified model corresponds to a regime of strong local co-localization where binding is effectively instantaneous, whereas the non-degenerate model captures a more realistic buffered regime in which reversible complex formation attenuates spatial feedback. The resulting differences in pattern amplitude and stability mirror known biophysical roles of LLPS in regulating metabolite gradients, enhancing local reaction efficiency, and preventing uncontrolled accumulation of intermediates.
Overall, this study establishes a quantitative link between enzymatic binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as LLPS. The framework developed here can be readily extended to more complex metabolic pathways, multi-enzyme assemblies, and other compartmentalized biochemical systems, offering a versatile foundation for understanding how molecular self-organization shapes cellular metabolism and morphogenesis.
8. Acknowledgments
The author warmly thanks Prof. Stepan Timr (J. Heyrovský Institute of Physical Chemistry, Czech Academy of Sciences) for his constant support, stimulating discussions, and many helpful suggestions that greatly improved the quality of this work. FF is also grateful for support from the National Group of Mathematical Physics (GNFM-INdAM).