### v2

# Enzyme-Substrate Complex Formation Modulates Diffusion-Driven Patterning in Metabolic Pathways

Abstract.
We investigate how enzymatic binding kinetics regulate diffusion-driven instabilities in a two-step metabolic pathway. Starting from a mechanistic description in which the substrate reversibly binds to the first enzyme before catalytic conversion, we formulate two reaction–diffusion models: a simplified system with effective kinetics and an extended model that explicitly includes the enzyme–substrate complex. The latter exhibits a structural degeneracy at critical parameter values due to a continuous family of homogeneous equilibria. To enable direct comparison and analytical progress, we introduce a reduced non-degenerate formulation via a quasi-equilibrium closure that encodes the influence of complex formation into effective reaction terms while preserving the nonlinear coupling between catalytic turnover and spatial transport.
We show that explicit enzyme–substrate binding shifts the homogeneous steady state, modifies relaxation dynamics, and substantially alters the size and location of the Turing instability region relative to the simplified model. Numerical simulations are in close agreement with weakly nonlinear predictions, illustrating how reversible binding reshapes pattern selection and slows the development of spatial heterogeneity.
These results establish a quantitative link between enzyme–substrate binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization, including structures associated with liquid–liquid phase separation (LLPS). The proposed framework provides a mechanistic route by which association, dissociation, and catalytic rates jointly regulate the robustness and structure of spatial metabolic patterns, and can be extended to broader classes of compartmentalized biochemical networks.
[…]
F. FARIVAR
[…]
1. Introduction
Understanding the complex interactions in biochemical reaction networks is essential for uncovering the principles of cellular function and metabolic regulation. Among the key features of these systems are reaction-diffusion processes, in which chemical species diffuse through space and interact in nonlinear ways. These interactions can give rise to striking spatial organizations, most notably Turing patterns, which are periodic structures that emerge spontaneously from homogeneous initial conditions when reaction kinetics and diffusion rates satisfy certain instability criteria. Since the pioneering work of Turing (Turing (1952)), reaction-diffusion mechanisms provide a robust mathematical framework for the generation of spatial patterns in biological systems (Maini et al. (2012); Maini and Woolley (2019)). ranging from chemical oscillators (Epstein et al. (2012), Gambino et al. (2015)) to developmental patterning in multicellular systems (Furusawa and Kaneko (1998); Giunta et al. (2021)).
Metabolic pathways in living organisms provide a particularly rich setting for studying such processes. These pathways consist of sequences of enzymatic reactions that convert substrates into intermediates and final products, often with multiple branches and feedback loops (Hafner and Hatzimanikatis (2021), Lee et al. (2025), Hill et al. (2015)). Spatial aspects of metabolism have received growing attention in recent years, as it has become clear that enzymes and metabolites are not always homogeneously distributed in the cytoplasm. Instead, cells frequently organize enzymes into clusters or into biomolecular condensates through liquid–liquid phase separation (LLPS) (Milicevic et al. (2022), Li et al. (2025), Chen et al. (2022)). Such mesoscale organization can accelerate metabolic fluxes, buffer fluctuations, and regulate competition at branch points (Bevilacqua et al. (2024), Lu et al. (2021), Lim and Clark (2024), Krainer et al. (2021); Liu et al. (2023)). Liquid‒liquid phase separation is a fundamental mechanism of sepsis (Chen et al. (2025)) From a theoretical perspective, reaction-diffusion models offer a natural framework to investigate how enzyme clustering and LLPS influence the spatial dynamics of metabolism (Castellana et al. (2014); Mukherjee et al. (2024), Buchner et al. (2013); Cieza Huaman (2022); Nakasone and Terazima (2021); Rigano (2022); Kim et al. (2024)). For instance, liquid-liquid phase separation (LLPS) on cell membranes has been described using reaction-diffusion models, which demonstrate that enzymatic feedback loops may drive out-of-equilibrium dynamics that replicate classical phase separation properties as nucleation, coarsening, and domain creation Rigano (2022). A straightforward reaction-diffusion framework called the swarm model has been put out to explain the molecular condensation processes that occur in cells. It demonstrates how these condensates emerge out of equilibrium and may be controlled by nucleation and cooperativity dynamics. It captures important aspects of membraneless organelle development, including nucleation, clustering, diffusion, and particle exchange (Cieza Huaman (2022)).
In this work, we introduce Models 1, 2, and 3 to investigate the conditions for Turing pattern formation in enzyme-mediated metabolic pathways. The Table 1 summarizing all parameters contained in Models 1 and 2 followed as:
Parameter Description (with units)
[…]
$\alpha_{0}$ Relaxation rate toward $C_{0}^{*}$   [s-1]
[…]
$k_{d_{1}}$ Dissociation rate of $E_{1}S_{0}$   [s-1]
$k_{cat}$ Catalytic turnover rate to $S_{1}$   [s-1]
$\beta$ Degradation rate of $C_{1}$   [s-1]
[…]
Table 1. Summary of parameters used in Models 1 and 2 with physical units.
Unless otherwise stated, all parameters appearing in the models and numerical simulations use the dimensional units listed in Table 1.
The novelty of this work lies in combining mechanistic enzyme–substrate kinetics with diffusion and cross-diffusion to study the emergence of spatial metabolic organization. We introduce a non-degenerated reaction–diffusion formulation that explicitly incorporates reversible enzyme–substrate complex formation and derive a reduced model via a quasi-equilibrium approximation that preserves the essential nonlinear feedback between catalysis and spatial transport. By comparing this model with a classical simplified enzymatic pathway, we demonstrate how complex formation shifts the homogeneous steady state, alters relaxation dynamics, and significantly modifies the Turing instability region. To our knowledge, this is the first analysis quantifying how enzymatic binding kinetics reshape diffusion-driven instabilities in two-step metabolic pathways, providing a mechanistic link between enzyme-mediated feedback and mesoscale organization such as liquid–liquid phase separation.
1.1. Model 1: The primary framework.
As a starting point, we consider a simplified two-step metabolic pathway (Castellana et al. (2014)): $S_{0}\xrightarrow{E_{1}}S_{1}\xrightarrow{E_{2}}P,$ where $S_{0}$, $S_{1}$, and $P$ denote substrate, intermediate, and final product, respectively. The substrate $S_{0}$ is converted to the intermediate $S_{1}$ by enzyme $E_{1}$, which in turn is converted to $P$ by enzyme $E_{2}$. Following the phenomenological formulation of (Castellana et al. (2014)), we model the spatio-temporal evolution of the substrate and intermediate as
$\begin{cases}\dfrac{\partial C_{0}}{\partial t}=-\alpha_{0}(C_{0}-C_{0}^{*})-k_{1}n_{1}C_{0}+D_{0}\nabla^{2}C_{0}-d\,C_{0}\nabla^{2}C_{1},\\[6.0pt] \dfrac{\partial C_{1}}{\partial t}=-k_{2}n_{2}C_{1}-\beta C_{1}+k_{1}n_{1}C_{0}+D_{1}\nabla^{2}C_{1}.\end{cases}$ (1.1)
Here $C_{0}$ and $C_{1}$ denote the concentrations of $S_{0}$ and $S_{1}$, while $n_{1}$ and $n_{2}$ represent fixed enzyme levels. The parameter $\alpha_{0}$ drives $C_{0}$ toward its homeostatic value $C_{0}^{*}$, $k_{1}n_{1}$ and $k_{2}n_{2}$ are effective catalytic and consumption rates, and $\beta$ is the degradation rate of $C_{1}$. For $\beta=0$, the intermediate can accumulate indefinitely, eliminating steady-state flux through the pathway.
[…]
This coupling does not depend on the gradient of $C_{1}$ but on its curvature: at locations where $C_{1}$ attains a local maximum, $\nabla^{2}C_{1}<0$ and therefore $d\,C_{0}\nabla^{2}C_{1}>0$, implying that $C_{0}$ increases at such sites. Conversely, $C_{0}$ decreases near local minima ($\nabla^{2}C_{1}>0$). Thus, substrate accumulates preferentially near regions where the intermediate forms concave peaks, producing an effective recruitment mechanism without invoking direct molecular attraction.
Such curvature-dependent transport is consistent with mesoscale biophysical processes. Curvature-sensing proteins in membranes preferentially bind to negatively curved regions (Arnold et al. (2025)), and curvature gradients can drive molecular fluxes in cells and tissues (Schamberger et al. (2023)). Analogous behavior occurs in LLPS condensates and metabolon-like assemblies, where spatial variations in viscosity, binding affinity, and chemical potential bias metabolite transport (Castellana et al. (2014); Zwicker et al. (2017)). The term $d\,C_{0}\nabla^{2}C_{1}$ therefore provides a minimal and biophysically motivated representation of curvature-driven substrate recruitment.
LLPS proceeds through nucleation, maturation, and functional biochemical activity (Hyman et al. (2014); Banani et al. (2017)). Model 1 corresponds to the latter two stages: we assume that a condensate has already formed and maintains approximately constant internal enzyme concentrations $n_{1}$ and $n_{2}$, while the metabolites $C_{0}$ and $C_{1}$ remain mobile and diffuse into, out of, and within the condensed phase. The curvature-driven term $d\,C_{0}\nabla^{2}C_{1}$ captures substrate recruitment into regions of high intermediate accumulation, a hallmark of the mature, functional stage of LLPS-mediated metabolic organization.
Model 1 thus serves as a coarse-grained baseline for studying generic reaction–diffusion feedbacks in enzyme-containing LLPS compartments and provides the foundation upon which mechanistic extensions (Model 2) can be built.
1.2. Model 2: Explicit enzyme–substrate interactions.
Model 1 captures basic spatial coupling but treats catalysis as an effective one-step process and does not resolve the transient enzyme–substrate complex. In cellular environments, however, enzymes bind substrates reversibly before turnover (Keener and Sneyd (2009)), and the mobility of such complexes—especially inside co-localized assemblies such as liquid–liquid phase-separated (LLPS) condensates—may differ substantially from that of free metabolites.
To account for these effects, we introduce an additional dynamical variable representing the enzyme–substrate complex $C_{E_{1}S_{0}}$. The underlying biochemical scheme,
$E_{1}+S_{0}\xrightleftharpoons[K_{d_{1}}]{k_{a_{1}}}E_{1}S_{0}\xrightarrow{k_{cat}}E_{1}+S_{1},$
motivates the extended reaction–diffusion system
$\begin{cases}\dfrac{\partial C_{0}}{\partial t}=-\alpha_{0}(C_{0}-C_{0}^{*})-k_{d_{1}}n_{1}\,C_{E_{1}S_{0}}+\gamma_{0}\nabla^{2}C_{0}-\gamma\,C_{0}\nabla^{2}C_{1},\\[6.0pt] \dfrac{\partial C_{E_{1}S_{0}}}{\partial t}=k_{a_{1}}n_{1}\,C_{0}-(k_{d_{1}}+k_{cat})C_{E_{1}S_{0}}+\gamma_{\mathrm{comp}}\nabla^{2}C_{E_{1}S_{0}},\\[6.0pt] \dfrac{\partial C_{1}}{\partial t}=-(k_{2}n_{2}+\beta)\,C_{1}+k_{cat}\,C_{E_{1}S_{0}}+\gamma_{1}\nabla^{2}C_{1},\end{cases}$ (1.2)
Remark. Equation (1.2) is not intended as a strict microscopic mass-action description. Instead, $C_{E_{1}S_{0}}$ acts as an effective complex-density field summarizing local co-localization of enzyme and substrate—for example within LLPS condensates or metabolon clusters. The terms $k_{a_{1}}n_{1}C_{0}$ and $k_{d_{1}}n_{1}C_{E_{1}S_{0}}$ therefore represent coarse-grained recruitment and release processes characteristic of mesoscale organization (Castellana et al. (2014); Zwicker et al. (2017)).
Model 2 extends the phenomenological dynamics of Model 1 by introducing an explicit intermediate while preserving the same basic mechanism: feedback between catalytic activity and spatial metabolite organization. Model 1 provides a minimal two-species description, whereas Model 2 offers a more mechanistic formulation linking enzyme binding, catalytic turnover, and condensate-mediated localization.
1.3. Model 3: Reduced non-degenerate formulation.
While explicitly including the enzyme–substrate complex enriches the model biologically, the homogeneous dynamics of Model 2 becomes degenerate at the critical relation $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$, where the system admits a one-dimensional manifold of equilibria and the Jacobian has a zero eigenvalue. This non-hyperbolicity obstructs the direct application of standard weakly nonlinear (WNL) theory in its simplest form, which typically assumes an isolated equilibrium at the bifurcation point.
To enable direct comparison with the primary two-variable pathway and to restore analytical tractability, we introduce a reduced two-variable non-degenerated model (Section 3.2) by eliminating $C_{E_{1}S_{0}}$ through a quasi-equilibrium closure that neglects its diffusive contribution at leading order and expresses the complex concentration in terms of the substrate. The resulting model preserves the mechanistic influence of complex formation through modified effective reaction coefficients, while retaining the same cross-diffusion structure that encodes spatial recruitment in LLPS-like environments.
Throughout this work we impose homogeneous Neumann (zero–flux) boundary conditions, for all models and all parameters are primary considered non-negative. These conditions ensure mass conservation and reflect the fact that metabolites and enzyme–substrate complexes do not cross the boundary of the system.
Although the numerical simulations presented in Section 6 are performed in one spatial dimension, the reaction–diffusion mechanism is not restricted to 1D. The same chemical interactions and diffusion laws apply in two and three dimensions, and the Turing conditions are dimension-independent: what changes is only the geometry of the domain and the admissible wave numbers. In higher dimensions, additional pattern morphologies (spots, stripes, labyrinths, hexagons)(Murray (2007), Gambino et al. (2013, 2019)) may occur, but the biochemical interpretation—enzymatic feedback coupled with substrate/intermediate mobility—remains unchanged. Thus, the 1D setting captures the essential instability mechanism while allowing clearer analytical and numerical comparison between the simplified and non-degenerated models.
The remainder of this paper is organized as follows. In Section 2, we analyze the local dynamics of the primary model and derive the conditions for Turing instability driven by diffusion and cross-diffusion effects. In Section 3, we examine the equilibrium structure of the extended three-variable system, showing that it is degenerate. To address this, we employ a quasi-equilibrium approximation (Keener and Sneyd (2009); Murray (2007)) to obtain a reduced non-degenerate model, for which we investigate both local stability and the onset of Turing instability. Section 4 explores how variations in reaction and diffusion parameters affect the extent of Turing regions in Models 1 and 3. In Section 5, we apply weakly nonlinear (WNL) analysis to the primary and reduced models to derive amplitude equations describing the nonlinear evolution of spatial patterns. These analytical predictions are then compared with numerical simulations in Section 6, confirming the validity of the theoretical results. Finally, Section 7 presents concluding remarks and discusses the broader implications of our findings for understanding spatial regulation in enzyme-mediated and phase-separated biochemical systems.
2. Simplified Two-Step Enzyme Model
In this section, we present linear analysis of the primary model (1.1) around the unique steady state of the system. For all positive non-negative parameters, this system has an unique coexistence steady state $E^{*}=(C_{0}^{E},C_{1}^{E})$ that is:
$C_{0}^{E}=\dfrac{\alpha_{0}C_{0}^{*}}{\alpha_{0}+k_{1}n_{1}},\;\;C_{1}^{E}=\dfrac{k_{1}n_{1}\alpha_{0}*C_{0}^{*}}{(k_{2}n_{2}+\beta)(\alpha_{0}+k_{1}n_{1})},$ (2.1)
Linear analysis of the kinetic around $E^{*}$ demonstrates that $E^{*}$ is always stable since
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-k_{1}n_{1}&&0\\ k_{1}n_{1}&&-\beta-k_{2}n_{2}\end{pmatrix}$, (2.2)
So the characteristic function is obtained for $|\lambda I-J(E^{*})|=\lambda^{2}+\lambda tr(J)+det(J)=0,$ that has two negative eigenvalues $\lambda_{1}=-\alpha_{0}-k_{1}n_{1},\;\;\lambda_{2}=-\beta-k_{2}n_{2}.$
Proposition 2.1.
The system (1.1) does not go under Hopf bifurcation since all parameters are non negative and $tr(J)>0$.
2.1. Turing Analysis
According to the Turing’s statement, a steady state can be unstable in presence of diffusion. In order to investigate Turing instabilities, one needs to linearize whole system at $E^{*}$, which provides:
$\mathcal{D}(E^{*})=\begin{pmatrix}D_{0}&&-dC_{0}^{E}\\ 0&&D_{1}\end{pmatrix}$, (2.3)
and for $W=[C_{0}-C_{0}^{E},C_{1}-C_{1}^{E}]$, we have
$\dot{W}=J(E^{*})W+\mathcal{D}(E^{*})\nabla^{2}W,$
In order to find the solution of the linearized model with replace the system with $W=e^{ikx+\lambda t}$, in which $k$ and $\lambda$ imply wave number and growth rate correspondingly. Therefore, dispersion relation of the model is determined as $|\lambda I-J+k^{2}D|=0$. Thus, the dispersion relation is given by:
$\lambda^{2}+A(k^{2})\lambda+B(k^{2})=0,$ (2.4)
$A(k^{2})=k^{2}(D_{1}+D_{0})+\alpha_{0}+k_{1}n_{1}+\beta+k_{2}n_{2},$ (2.5)
According to the Turing analysis, $E^{*}$ is unstable as one eigenvalue of the dispersion relation is negative, that requires $B(k^{2})<0$ for a range of wave numbers $k_{a},k_{b}]$, where
$B(k^{2})$ $=k^{4}D_{0}D_{1}+k^{2}[-D_{0}(-\beta-k_{2}n_{2})-D_{1}(-\alpha_{0}-k_{1}n_{1})-k_{1}n_{1}dC0_{E}^{*}]$ (2.6)
$+(-k_{1}n_{1}-\alpha_{0})(-\beta-k_{2}n_{2})=b_{1}k^{4}+b_{2}k^{2}+b_{3},$ (2.7)
Turing necessary condition obtains that in the Turing threshold
$B(k_{min}^{2})=0,$ (2.8)
which gives
$k^{2}_{min}=\dfrac{-b_{2}}{2b_{1}},$ (2.9)
Since only positive wave numbers are physically meaningful and $b_{1}>0$, therefore, $b_{2}$ must be negative. This provides that critical Turing bifurcation parameter is $d$ and it is determined by:
$d>\dfrac{-D_{0}(-k_{2}n_{2}-\beta)-D_{1}(-k_{1}n_{1}-\alpha_{0})}{k_{1}n_{1}C0_{E}^{*}},$ (2.10)
So to find the critical wave number we replace $d_{c}=\eta_{1}/\eta_{2}+\varepsilon$ where
$\eta_{1}=-D_{0}(-k_{2}n_{2}-\beta)-D_{1}(-k_{1}n_{1}-\alpha_{0}),$ (2.11)
$\eta_{2}=k_{1}n_{1}C0_{E}^{*},$ (2.12)
into (2.8) which obtains that
$\varepsilon=\dfrac{2\sqrt{D_{0}D_{1}(-\alpha_{0}-k_{1}n_{1})(-\beta-k_{2}n_{2})}}{k_{1}n_{1}C0_{E}^{*}},$ (2.13)
and consequently
$k_{min}^{2}=\sqrt{\dfrac{(-\alpha_{0}-k_{1}n_{1})(-\beta-k_{2}n_{2})}{D_{0}D_{1}}}.$ (2.14)
Hence, the system goes to the pattern formation if necessary condition (2.10) is satisfied.
Remark 2.1.
It is easy to check that in absence of cross-diffusion $(d=0)$, they system remain stable since $b_{2}>0$ for any given parameters. So, cross-diffusion term is only key mechanism of emergence of Turing patterns.
3. Modified Model and Quasi-Equilibrium Approximation
3.1. Equilibria and Stability Analysis of the Modified Model
To obtain a more mechanistic representation of the enzymatic conversion process, the reaction scheme is extended to include the reversible formation of the enzyme–substrate complex and its catalytic conversion into the intermediate. This leads to the three–species reaction–diffusion system (1.2), which we first analyse in the spatially homogeneous setting.
Equilibria. Setting time derivatives to zero in the homogeneous system yields two types of steady states:
  • (1)
    The equilibrium $E_{1}^{*}=(C_{0}^{*},0,0),$ which exists whenever $k_{a_{1}}n_{1}-(\,k_{d_{1}}+k_{cat}\,)\neq 0$.
  • (2)
    When the condition $k_{a_{1}}n_{1}-(\,k_{d_{1}}+k_{cat}\,)=0$ holds, the steady–state equations become linearly dependent and admit a one–parameter family of equilibria $E_{2}^{*}(\xi)=\left(\frac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}\,\xi}{\alpha_{0}},\;\xi,\;\frac{k_{cat}}{k_{2}n_{2}+\beta_{2}}\,\xi\right),\qquad\xi\in\mathcal{I},$ where $\mathcal{I}$ is chosen to ensure nonnegativity of the components. This degeneracy corresponds to a continuous curve of equilibria.
Stability of $E_{1}^{*}$. Linearization of system (1.2) around $E_{1}^{*}$ gives
$J(E_{1}^{*})=\begin{bmatrix}-\alpha_{0}&-k_{d_{1}}n_{1}&0\\[2.84526pt] 0&k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})&0\\[2.84526pt] 0&k_{cat}&-(k_{2}n_{2}+\beta_{2})\end{bmatrix}$, (3.1)
whose eigenvalues are $\lambda_{1}=-\alpha_{0},\qquad\lambda_{2}=-(k_{2}n_{2}+\beta_{2}),\qquad\lambda_{3}=k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat}).$ Because $\alpha_{0}>0$ and $(k_{2}n_{2}+\beta_{2})>0$, the first two eigenvalues are strictly negative, and the sign of $\lambda_{3}$ determines the local stability of $E_{1}^{*}$.
Theorem 3.1.
Let $E_{1}^{*}$ denote the equilibrium of system (1.2). Then:
  1. (1)
    If $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})<0$, then $E_{1}^{*}$ is locally asymptotically stable.
  2. (2)
    If $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})>0$, then $E_{1}^{*}$ is unstable.
Proof.
Immediate from the eigenvalues of $J(E_{1}^{*})$, as described above. ∎
Degenerate case: line of equilibria. At the critical value $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0,$ the Jacobian at any equilibrium point $E_{2}^{*}(\xi)$ possesses two strictly negative eigenvalues and one zero eigenvalue. Thus every point on the equilibrium curve $\mathcal{M}=\{E_{2}^{*}(\xi)\}$ is non–hyperbolic: perturbations orthogonal to $\mathcal{M}$ decay exponentially, while perturbations tangent to $\mathcal{M}$ neither grow nor decay. This neutral direction reflects the structural degeneracy of the steady–state equations and the presence of a continuous family of equilibria.
Geometrically, the set $\mathcal{M}$ forms a one–dimensional invariant manifold of steady states, with exponential attraction in the two transverse directions. Perturbations therefore relax rapidly onto the neighbourhood of $\mathcal{M}$, after which the dynamics remains close to this manifold.
3.2. Equilibrium Analysis and Derivation of the Non-degenerated Model
To obtain the equilibrium concentrations $C_{0}^{*}$, $C_{1}^{*}$, and $C_{E_{1}S_{0}}^{*}$, we first rewrite the steady-state conditions corresponding to system (1.2). From the first equation we obtain the linear relation $C_{0}=C_{0}^{*}-\frac{k_{d_{1}}n_{1}}{\alpha_{0}}\,C_{E_{1}S_{0}}.$ Under the simplifying condition $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$, the second equation becomes singular and the homogeneous kinetic system is degenerate along a one-dimensional manifold of equilibria. Solving the third equation for the intermediate $C_{1}$ yields $C_{1}=\frac{k_{cat}}{k_{2}n_{2}+\beta}\,C_{E_{1}S_{0}}.$
Because of this degeneracy, it is useful to construct a reduced model in which the enzyme–substrate complex is expressed in terms of the substrate concentration. To this end, we adopt a quasi-equilibrium approximation for $C_{E_{1}S_{0}}$ by assuming that, on the spatial and temporal scales of interest, its diffusive contribution is small compared to the local reaction term. Neglecting the term $D\nabla^{2}C_{E_{1}S_{0}}$ in the second equation of (1.2) then gives the algebraic approximation $(k_{d_{1}}+k_{cat})\,C_{E_{1}S_{0}}\approx k_{a_{1}}n_{1}\,C_{0},$ and thus
$C_{E_{1}S_{0}}\approx\frac{k_{a_{1}}n_{1}}{k_{d_{1}}+k_{cat}}\,C_{0}.$ (3.2)
Substituting (3.2) into the remaining equations of system (1.2) yields the following reduced, or non-degenerated, two-variable model:
$\begin{cases}\dfrac{\partial u_{0}}{\partial t}=-\alpha_{0}\big(u_{0}-u_{in}^{*}\big)-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}}\,u_{0}+\gamma_{0}\nabla^{2}u_{0}-\gamma\,u_{0}\,\nabla^{2}u_{1},\\[5.69054pt] \dfrac{\partial u_{1}}{\partial t}=-(k_{2}n_{2}+\beta)\,u_{1}+\dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}\,u_{0}+\gamma_{1}\nabla^{2}u_{1}.\end{cases}$ (3.3)
The non-degenerated system (3.3) has the same structure as the primary model (1.1), but with modified effective coefficients that incorporate the contribution of the enzyme–substrate complex via the quasi-equilibrium approximation (3.2). In this formulation, $C_{E_{1}S_{0}}$ no longer appears as an independent variable, while the mechanistic effects of binding and catalytic conversion are retained through the renormalized reaction terms. This reduced model provides a convenient comparison to the primary pathway without altering the full three-species dynamics used for the pattern-formation analysis.
3.3. Linear Analysis of the Non-degenerated Reduced Model
The reduced non-degenerated system (3.3), obtained via the quasi-equilibrium approximation of the enzyme-substrate complex, admits a unique positive coexistence steady state given by
$u_{0}^{*}=\dfrac{\alpha_{0}u_{in}^{*}(k_{d_{1}}+k_{cat})}{\alpha_{0}(k_{d_{1}}+k_{cat})+k_{a_{1}}k_{d_{1}}n_{1}^{2}},\qquad u_{1}^{*}=\dfrac{k_{a_{1}}k_{cat}n_{1}\alpha_{0}u_{in}^{*}}{(k_{d_{1}}+k_{cat})(k_{a_{1}}k_{d_{1}}n_{1}^{2})}.$ (3.4)
All parameters are assumed to be positive.
This model contains all positive parameters. Linearizing the model at $E^{*}=(u_{0}^{*},u_{1}^{*})$ gives
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n^{2}}{k_{d_{1}}+k_{cat}}&&0\\ \dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}&&-\beta-k_{2}n_{2}\end{pmatrix},\;\;\;\mathcal{D}(E^{*})=\begin{pmatrix}\gamma_{0}&&-\gamma u_{0}^{*}\\ 0&&\gamma_{1}\end{pmatrix}$. (3.5)
This system is locally stable at $E^{*}$ simultaneously. And it admits Turing instability in presence of diffusion with necessary condition
$\gamma>\dfrac{-\gamma_{0}(-k_{2}n_{2}-\beta)-\gamma_{1}(-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}})}{\dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}u_{0}^{*}},$ (3.6)
So $\gamma$ is Turing bifurcation parameter. Moreover, critical wave number is respectively determined
$k_{min}^{2}=\sqrt{\dfrac{(\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n_{1}^{2}}{k_{d_{1}}+k_{cat}})(-k_{2}n_{2}-\beta)}{\gamma_{0}\gamma_{1}}},$
Figure 1. Phase-plane comparison of the simplified and non-degenerated enzyme-pathway models. Arrows represent vector fields (dark gray: primary model; light gray: non-degenerated model). Red/blue and magenta/green curves denote the corresponding nullclines, whose intersections yield homogeneous steady states (dots). Trajectories from identical initial conditions are shown in magenta and green. The non-degenerated formulation shifts the steady state toward higher intermediate concentration and exhibits slower relaxation, consistent with nonlinear feedback between catalysis and compartmentalization.
We have investigated numerically necessary conditions of appearance of Turing instabilities of both models (1.1) and (3.3) of a given set of parameters (see (Figure 2)). In this figure, plot of $B(k^{2})$ demonstrate as bifurcation parameters $d$ of (1.1) and $\gamma$ of (3.3) crosses their critical values, necessary conditions (2.10) and (3.6) are satisfied and correspondingly for both models $B(k^{2})$ admit negative value and respectively $\lambda(k^{2})$ admit positive values, which states appearance of pattern formation.
Figure 2. Plots of polynomials $B(k^{2})$ of two models (3.3) (red plots) and (1.1) (black plots) (right figure). Plots of dispersion relation polynomials $\lambda^{2}(k^{2})$ (left figure). Given parameters are $\alpha_{0}=0.01(s^{-1})$, $k_{1}=0.05(\mu M^{-1}s^{-1})$, $k_{2}=0.02(\mu M^{-1}s^{-1})$, $k_{d_{1}}=0.01s^{-1}$, $k_{a_{1}}=0.03(\mu M^{-1}s^{-1})$, $k_{cat}=0.1s^{-1}$, $\beta=0.01(s^{-1})$, $C_{0}^{*}=0.1(\mu Ms^{-1})$, $n_{1}=10(\mu M)$, $n_{2}=5(\mu M)$, $\gamma_{0}=0.1(\mu m^{2}s^{-1})$, $\gamma_{1}=0.01(\mu m^{2}s^{-1})$. where in this figures $d_{c}$ and $\gamma_{c}$ are Turing bifurcation parameter of model (1.1) and (3.3) respectively.
A qualitative comparison between the kinetic dynamics of the simplified model (1.1) and the reduced non-degenerated model (3.3) is shown in Figure 1. The phase–plane analysis highlights several key differences: (i) the nullclines of the non-degenerated model intersect at a steady state with higher intermediate concentration, (ii) trajectories relax more slowly toward equilibrium, reflecting the additional nonlinear feedback introduced by enzyme–substrate binding, and (iii) the vector fields of the two systems display distinct curvature, indicating that the mechanistic correction introduced by the approximation reduction significantly alters the underlying kinetics.
4. Turing region
To understand how parameters influence the emergence of Turing instability in the full enzymatic model, and to compare these effects with those predicted by the reduced non-degenerated system, we investigate the corresponding Turing regions in parameter space. In order to achieve this, we need to consider positivity conditions of both $E^{*}$, stability conditions in absence of diffusion. According to the (2.1) and (3.4) for the parameter ranges considered in this study, the coexistence steady states of both models are positive and locally asymptotically stable in the absence of diffusion. Thus, there are only Turing conditions (2.10), and (3.6). According to these conditions, we investigate variation of parameters in plane $(n_{1},d),\;(n_{1},\gamma)$. Figure 2, demonstrates that for the given data set, the bifurcation parameters of models (1.1) and (3.3) are $d_{c}=2.7190$ and $\gamma_{c}=5.4035$ correspondingly.
4.0.1. Influence of Intermediate Diffusion on the Stability of Enzyme–Substrate Spatial Organization in model (1.1)
To investigate the influence of diffusion on pattern formation and aggregation, we examined how variations in the diffusion coefficients of both the substrate ($C_{0}$) and the intermediate ($C_{1}$) affect the emergence of spatial structures. Our analysis reveals that increasing either $D_{0}$ or $D_{1}$ leads to a progressive reduction of the Turing instability region. As diffusion becomes more efficient, the concentration gradients of both chemical species are rapidly smoothed out, preventing the formation of localized substrate-enriched zones. Consequently, the spatial structures (lobes or peaks) that would otherwise arise from enzyme–substrate feedback are flattened, and the system tends toward a spatially homogeneous steady state. This stabilizing effect of molecular mobility is clearly visible in Figure 3, where the Turing domain diminishes with increasing $D_{0}$ and $D_{1}$.
Figure 3. Variation of Turing regions in plane $(n_{1},d)$ as the parameters $D_{0}$ (left plot) and $D_{1}$ (right plot) increase. Red dots in the plots indicate $d_{c}$ threshold for each given $D_{0}$ or $D_{1}$. Parameters have chosen like Figure 2.
4.0.2. Influence of Intermediate Diffusion on the Stability of Enzyme–Substrate Spatial Organization in model (3.3)
The analysis of Model (3.3) reveals that the diffusion of the intermediate plays a crucial role in regulating the onset of spatial organization. In Figure 4, as the diffusivity of the intermediate increases, the Turing region progressively shrinks and the instability threshold $\gamma_{c}(n_{1})$ shifts toward higher values of the cross-diffusion parameter. This behavior indicates that enhanced intermediate mobility tends to equalize its spatial distribution, thereby weakening the gradients necessary to sustain localized substrate–intermediate interactions. From a biochemical perspective, this corresponds to a reduction in spatial compartmentalization: the mobile intermediate rapidly redistributes throughout the domain, preventing the formation of enzyme–substrate-enriched zones and driving the system toward a spatially homogeneous steady state.
In both the primary and non-degenerated models, increasing the substrate diffusivity $D_{i},\gamma_{i}$ reduces the Turing region, indicating that higher substrate mobility stabilizes the homogeneous state. However, in the non-degenerated model, where the cross-diffusion coupling depends explicitly on substrate concentration and reaction kinetics, this stabilizing influence is stronger. Thus, while both models exhibit diffusion-mediated suppression of pattern formation, the mechanistic coupling in the non-degenerated model amplifies the homogenizing effect.
Figure 4. Variation of Turing regions in plane $(n_{1},\gamma)$ as the parameters $\gamma_{0}$ (left plot) and $\gamma_{1}$ (right plot) increase. Red dots in the plots indicate $\gamma_{c}$ threshold for each given $\gamma_{0}$ or $\gamma_{1}$. Parameters have chosen like Figure 2.
4.0.3. Influence of Kinetic Parameters on the Stability of Enzyme–Substrate Spatial Organization in Model (3.3)
Figures 5 illustrate the influence of the kinetic parameters $k_{a_{1}}$, $k_{d_{1}}$, and $k_{cat}$ on the onset of diffusion-driven instability. In all cases, the system displays the characteristic U-shaped Turing boundary, but the extent and position of the unstable region are strongly affected by the reaction rates. Increasing the association rate $k_{a_{1}}$ enlarges the Turing region and shifts it toward higher $n_{1}$, indicating that stronger substrate–enzyme binding enhances the nonlinear coupling responsible for spatial self-organization. Conversely, increasing the dissociation rate $k_{d_{1}}$ diminishes the Turing domain, showing that faster complex dissociation stabilizes the homogeneous steady state. A similar expansion of the Turing region is observed for higher catalytic rates $k_{cat}$, where a more efficient turnover of the intermediate amplifies local concentration gradients and reinforces the enzyme–substrate feedback. Overall, strong association and high catalytic activity favor the emergence of spatial structures, whereas rapid dissociation weakens the coupling and suppresses pattern formation.
Figure 5. Variation of Turing regions in plane $(n_{1},\gamma)$ as the parameter $k_{cat},\;k_{d_{1}},\;k_{a_{1}}$ increases res. Parameters have chosen like Figure 2.
Figure 6. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left plot) as the parameter $\beta$ increases. Parameters have chosen like Figure 2.
Figure 7. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left) as the parameter $\alpha$ increases. Parameters have chosen like Figure 2.
The figures 6 and 7 investigate how Turing instability regions shift under variations of key parameters, namely $\beta$ and $\alpha_{0}$, in the $(n_{1},d)$ and $(n_{1},\gamma)$ planes. In both the $(n_{1},d)$ and $(n_{1},\gamma)$ plots, increasing $\beta$ leads to higher critical values for $d$ and $\gamma$, respectively. This indicates that larger degradation or reaction rates of the inhibitor suppress pattern formation, requiring stronger diffusion or cross-diffusion to induce Turing instability. Conversely, the $(n_{1},d)$ and $(n_{1},\gamma)$ plots with varying $\alpha_{0}$ show that decreasing $\alpha_{0}$ significantly broadens the Turing region. In $(n_{1},d)$ plane, when $\alpha_{0}\rightarrow 0$, the threshold for $d$ approaches zero, suggesting that slower decay of the activator greatly facilitates the onset of instability.
In summary, the figures illustrate that the Turing space is highly sensitive to kinetic parameters. Smaller values of $\beta$ and $\alpha_{0}$ enlarge the admissible region in the parameter space where spatial patterns can form. This implies that systems with slower degradation or decay kinetics are more susceptible to Turing-type pattern formation. These insights are particularly relevant for enzymatic and reaction-diffusion models in biological systems, where parameter tuning can drive or suppress morphogenetic processes.
Figure 8. Variation of Turing regions in plane $(n_{1},d)$ (right plot) and $(n_{1},\gamma)$ (left plot). Variation of Turing regions in plane $(n_{2},d)$ and $(n_{2},\gamma)$ (right plot). Parameters have chosen like Figure 2.
The plotted critical curves Figure 8 highlight how the onset of pattern formation depends nonlinearly on enzyme levels, with broader instability regions for low-to-moderate $n_{1}$, and increasingly suppressed Turing regions with high $n_{2}$. We analyze the critical diffusion parameters $d_{c}$ and $\gamma_{c}$ associated with Turing instability in two enzyme-based reaction-diffusion models. In the $n_{1}$-parameter space, both curves exhibit a non-monotonic, U-shaped behavior. This indicates that intermediate values of $n_{1}$ (the enzyme responsible for the first reaction step) are most conducive to pattern formation, requiring minimal diffusion or cross-diffusion to trigger instability. At both low and high $n_{1}$, the system becomes more stable, and higher values of $d_{c}$ and $\gamma_{c}$ are needed to induce patterns.
In contrast, the plots against $n_{2}$ show that both $d_{c}$ and $\gamma_{c}$ increase monotonically with enzyme concentration. This implies that the second enzyme, involved in downstream processes, stabilizes the system by diminishing feedback necessary for pattern formation. The widening gap between the two curves also suggests that the effect of cross-diffusion becomes increasingly significant as $n_{2}$ increases, reinforcing the role of downstream reactions in regulating spatial instabilities.
5. Weakly nonlinear Analysis of models
In this section we perform a weakly nonlinear analysis for the model (1.1) to predict the amplitude of the pattern near the Turing threshold. The weakly nonlinear analysis is based on the multiple scale methods (Gambino et al. (2012)). The similar strategy has applied on the model (3.3). Since near to the bifurcation the amplitude of the pattern (diffusion-driven instability) has slow temporal scale, then a new temporal scale is defined.
The solution of the original system is written as a weakly nonlinear expansion in the small control parameter $\varepsilon$. We choose $\varepsilon^{2}=\dfrac{d-d_{c}}{d_{c}}.$
The slow scale is obtained from the linear analysis: it is easy to prove that $\lambda\sim\varepsilon^{2}$ and, since the growth rate of the perturbation is proportional to the $exp(\lambda t)$, the slow time scale $T$ is order $\varepsilon^{2}$. Therefore, close to the threshold we separate the fast time $t$ and slow time $T=\varepsilon^{2}t$, so that time derivative is obtained as $\partial_{t}\rightarrow\partial_{t}+\varepsilon^{2}\partial_{T}$.
We separate the linear part from the nonlinear part:
$\partial_{t}\bm{w}=\mathcal{L}^{d_{c}}\bm{w}+\mathcal{Q}_{D}^{d_{c}}(\bm{w},\bm{\nabla^{2}w}),\qquad\bm{w}=[C_{0}-C_{0}^{E},C_{1}-C_{1}^{E}],$ (5.1)
and linear operator is defined as $\mathcal{L}^{d_{c}}=\Gamma J+\mathcal{D}^{d_{c}}\nabla^{2},$ where $\mathcal{D}$ and J{J} have defined in (2.2 and 2.3) and nonlinear operators $\mathcal{Q}_{D}^{d_{c}}(\bm{x},\bm{y})$ are introduced as: $\bm{x}=(x_{C_{0}},x_{C_{1}}),\;\bm{y}=(y_{C_{0}},y_{C_{1}})$,
$\mathcal{Q}_{D}(\bm{x},\bm{y})=\begin{bmatrix}-dx_{c_{0}}y_{c_{1}}\\ 0\end{bmatrix}$, (5.2)
And moreover the bifurcation parameter and the solution are expanded asymptotically
$d=d_{c}+\varepsilon^{2}d^{(2)}+O(\varepsilon^{4}),$ (5.3)
$\bm{w}=\varepsilon\bm{w}_{1}+\varepsilon^{2}\bm{w}_{2}+\varepsilon^{3}\bm{w}_{3}+O(\varepsilon^{4}),$ (5.4)
Therefore, the diffusion matrix is given in terms of perturbation parameters as:
$\mathcal{D}=\begin{bmatrix}D_{0}&-d_{c}C_{0}^{E}\\ 0&D_{1}\end{bmatrix}+\varepsilon^{2}\begin{bmatrix}0&-d^{(2)}C_{0}^{E}\\ 0&0\end{bmatrix}+O(\varepsilon^{4})$ (5.5)
and consequently
$\mathcal{Q}_{D}(\bm{w},\nabla^{2}\bm{w})$ $=\varepsilon^{2}d_{c}\mathcal{Q}_{D}(\bm{w}_{1},\nabla^{2}\bm{w}_{1})+\varepsilon^{3}d_{c}\{\mathcal{Q}_{D}(\bm{w}_{1},\nabla^{2}\bm{w}_{2})$ (5.6)
$+\mathcal{Q}_{D}(\bm{w}_{2},\nabla^{2}\bm{w}_{1})\}+\varepsilon^{4}d^{(2)}\mathcal{Q}_{D}(\bm{w}_{2},\nabla^{2}\bm{w}_{2})+O(\varepsilon^{5}),$ (5.7)
Now we replace all expansion in (5.3), (5.5),(5.4), and (5.6) and sort according to the order of $\varepsilon$. At $O(\varepsilon)$,we obtain the following linear problem:
$\mathcal{L}^{d_{c}}\bm{w}_{1}=0,\quad\bm{w}_{1}=A(T){\rho}\cos(k_{c}x),$ (5.8)
such that satisfying the homogeneous boundary conditions where $\rho$} belongs to the $ker(J-k^{2}_{c}\mathcal{D}^{d_{c}})$. In this stage the $A(T)$, the amplitude of the pattern is arbitrary and the vector $\rho$ is considered constant such whose normalization is
${\rho}=\begin{bmatrix}1\\ \rho_{2}\end{bmatrix}$, (5.9)
where $\rho_{2}=-\dfrac{-\alpha_{0}-k_{1}n_{1}-k_{c}^{2}D_{0}}{dC_{0}^{E}k_{c}^{2}}$.
Moreover, at $O(\varepsilon^{2})$ there is this linear equation which must be solved:
$\mathcal{L}^{d_{c}}\textbf{w}_{2}=\bm{F},$
According to the Fredholm alternative theorem, this equation has a solution if and only if $\langle\bm{F},\psi\rangle=0,$ where $\psi^{*}$ is defined at
$\psi=\begin{bmatrix}1\\ R_{1}^{*}\end{bmatrix}\cos(k_{c}x)=\psi^{*}\cos(k_{c}x),$
and $\langle.,.\rangle$ implied the scalar product in $\bm{L}^{2}(0,\dfrac{2\pi}{k^{c}})$ and $\psi^{*}\in Ker(J-k^{2}_{c}D^{d_{c}})^{\dagger}$ where $\dagger$ shows transpose of complex conjugate of the matrix.
In particular, $\bm{F}=-\dfrac{1}{4}A^{2}\sum_{i=0,2}\mathcal{M}_{i}(\rho,(\rho)\cos(ik_{c}x)$, in which $\mathcal{M}_{i}(\rho,\rho)=-i^{2}k_{c}^{2}\mathcal{Q}_{D}^{d_{c}}(\rho,\rho)$. Hence, the vector $\bm{w}_{2}$ is defined as
$\bm{w}_{2}=A^{2}(\bm{w}_{20}+\bm{w}_{22}\cos(2k_{c}x))$ so that $\mathcal{L}_{i}^{d_{c}}\bm{w}_{2i}=-\dfrac{1}{4}\mathcal{M}_{i}(\rho,\rho),\;i={0,2}$ and $\mathcal{L}_{i}=-i^{2}k^{2}_{c}D^{d_{c}}$.
In following, at $O(\varepsilon^{3})$ we obtain the linear problem
$\mathcal{L}^{d_{c}}\bm{w}_{3}=\bm{G},$ (5.10)
where $\bm{G}=\left(\dfrac{dA}{dT}\rho+A\bm{G}_{1}^{(1)}+A^{3}\bm{G}_{1}^{(3)}\right)\cos(k_{c}x)+A^{3}\bm{G}_{3}\cos(3k_{c}x),$ in which
$\bm{G}_{1}^{(1)}=k^{2}_{c}\begin{bmatrix}0&d^{(2)}C_{0}^{E}\\ 0&0\end{bmatrix}\rho,$
$\bm{G}_{1}^{(3)}=-k_{c}^{2}\mathcal{Q}_{D}(\rho,\bm{w}_{22})-\dfrac{1}{2}k_{c}^{2}\mathcal{Q}_{D}(\bm{w}_{22},\rho),$
$\bm{G}_{3}=3k^{2}_{c}\mathcal{Q}_{D}^{d_{c}}(\rho,\bm{w}_{22}),$
Finally, by applying the solvability condition, we obtain the Stuart-Landau equation for the amplitude
$\dfrac{\partial A}{\partial T}=\sigma A-LA^{3},$
In addition, solvability of the equation (5.10) depends on $\langle\bm{G},\psi\rangle=0$, whose coefficients are given by
$\sigma=-\dfrac{\langle\bm{G}_{1}^{(1)},\psi\rangle}{\langle\rho,\psi\rangle},\;\;\bm{L}=\dfrac{\langle\bm{G}_{1}^{(3)},\psi\rangle}{\langle\rho,\psi\rangle},$
In the Stuart-Landau equation the coefficient $\sigma$ is always positive while L could be either negative or positive, corresponding to a subcritical or supercritical bifurcation.
The nontrivial solution of the amplitude equation is
$\bm{A}_{\infty}=\sqrt{\dfrac{\sigma}{L}},$ (5.11)
which requires $L>0$, and therefore the result of this analysis is valid only for supercritical bifurcations.
Therefore, the asymptotic behavior of the solution is given by weakly nonlinear analysis of $O(\varepsilon^{3})$ is:
$\bm{W}=\varepsilon A\rho cos(k_{c}x)+\varepsilon^{2}A^{2}[\bm{w}_{20}+\bm{w}_{22}cos(2k_{c}x)]+O(\varepsilon^{3}).$ (5.12)
6. Numerical Results
In this section, we exhibit numerical results. Simulations were performed using the dimensional parameters reported in Table 1. The numerical results clearly show that the primary model (Model 1) and the non-degenerated model (Model 3) exhibit qualitatively similar Turing patterns, but with important quantitative differences in both amplitude and relaxation dynamics. In Model 1, the substrate–intermediate feedback is direct: the production of the intermediate $C_{1}$ depends instantaneously on the local availability of substrate $C_{0}$, and the curvature-driven cross-diffusion term $dC_{0}\nabla^{2}C_{1}$ reinforces this coupling by transporting substrate toward regions where $C_{1}$ accumulates. As a result, the effective positive feedback loop between $C_{0}$ and $C_{1}$ is strong, leading to patterns with relatively large amplitude and fast convergence toward the stationary profile (Figure 9).
Figure 9. Turing Pattern for given parameters: $n_{1}=3.5,\;n_{2}=1,\;k_{1}=k_{2}=2,\;\alpha_{0}=1,\;C_{0}^{*}=7,\;\beta=3,\;D_{0}=2,\;D_{1}=2.5$ that provides $C_{0}^{E}=0.87,\;C_{1}^{E}=1.22,\;k_{c}=2.8,\;d_{c}=9.51$.
Figure 10. Turing Pattern for given parameters: $n_{1}=3.5,\;n_{2}=1,\;k_{a_{1}}=5,\;k_{cat}=8,\;k_{d_{1}}=2.5;k_{2}=2,\;\alpha_{0}=1,\;C_{0}^{*}=7,\;\beta=3,\;D_{0}=2,\;D1=2.5$ that provides $u_{0}^{E}=0.44,\;u_{1}^{E}=1.19,\;k_{c}=2.5,\;d_{c}=14.76$.
In contrast, the non-degenerated model incorporates an explicit enzyme–substrate interaction, in which $S_{0}$ must first bind to the enzyme $E_{1}$ to form the complex $E_{1}S_{0}$ before being converted into the intermediate $S_{1}$. This reversible binding acts as a biochemical buffer: part of the substrate is temporarily stored in the complex, and the effective rate at which $S_{0}$ contributes to the production of $S_{1}$ is reduced to $\frac{k_{a_{1}}k_{\mathrm{cat}}n_{1}}{k_{d_{1}}+k_{\mathrm{cat}}}\;<\;k_{1}n_{1}.$ Consequently, the feedback between substrate and intermediate is weaker, and perturbations grow more slowly. This mechanistic buffering is visible in both the phase–plane trajectories and the numerical simulations: patterns in the non-degenerated model have smaller amplitude, smoother spatial profiles, and longer relaxation times (Figure 10).
Furthermore, the modified cross-diffusion structure of Model 3, where the coupling depends explicitly on the substrate concentration $u_{0}$, reduces the sensitivity of the system to spatial gradients. This further stabilizes the homogeneous state and explains why the Turing region of Model 3 is narrower and shifted toward higher diffusion or cross-diffusion values compared with Model 1.
Figure 11. Numerical results vs WNL of model 1 for given parameters: as Figure 9 and $\varepsilon^{2}=0.003$.
Figure 12. Numerical results vs WNL of model 2 for given parameters: as Figure 10 and $\varepsilon^{2}=0.0001$.
The excellent agreement between numerical simulations and the weakly nonlinear predictions (Figures 11-12) confirms that these differences arise from fundamental kinetic mechanisms rather than numerical artifacts. Overall, the numerical results demonstrate that explicit enzyme–substrate binding attenuates the strength of the feedback necessary for Turing instability, leading to weaker and more finely regulated spatial patterns. Biologically, this means that enzymatic complex formation acts as a natural buffering process that prevents excessive accumulation of intermediates and ensures more stable metabolic organization.
7. Conclusion
This work examined how enzymatic reaction kinetics and diffusion-driven transport interact to generate spatial metabolic organization. Linear stability and Turing analyses showed that diffusion and curvature-driven cross-diffusion can destabilize otherwise stable homogeneous steady states, producing spatial patterns through local enzyme–substrate feedback. Incorporating explicit enzyme–substrate complex formation introduces additional kinetic structure that modifies reaction balance and effective transport, thereby shifting the Turing threshold and altering the size and location of the instability region. The resulting reduced non-degenerate formulation removes the structural degeneracy associated with a manifold of equilibria and enables a complete weakly nonlinear (WNL) analysis while preserving the essential nonlinear coupling between catalysis and spatial organization.
Theoretical predictions from the WNL framework are strongly supported by numerical simulations, which show excellent agreement in both pattern amplitude and spatial profile. Comparison between the simplified and non-degenerate models reveals systematic kinetic and spatial differences: explicit enzyme–substrate binding shifts the homogeneous equilibrium toward higher intermediate concentrations, slows relaxation toward steady state, and leads to spatial patterns of reduced amplitude. These results highlight the buffering role of reversible complex formation and demonstrate how enzymatic binding kinetics regulate the strength of the feedback loop driving diffusion-driven instability.
Beyond their mathematical implications, these findings admit a natural interpretation in the context of liquid–liquid phase separation (LLPS). Within biomolecular condensates, enzymes and metabolites experience altered mobility, molecular crowding, and transient binding interactions that jointly modulate catalytic flux and spatial organization. The simplified model captures an effective, coarse-grained description of enzymatic feedback, whereas the non-degenerate formulation reflects a buffered regime in which reversible complex formation attenuates spatial feedback. The resulting differences in pattern amplitude and stability mirror known biophysical roles of LLPS in regulating metabolite gradients, enhancing reaction efficiency, and preventing uncontrolled accumulation of intermediates.
Overall, this study establishes a quantitative link between enzymatic binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as LLPS. The framework developed here can be extended to more complex metabolic pathways, multi-enzyme assemblies, and other compartmentalized biochemical systems, providing a versatile foundation for understanding how molecular self-organization shapes cellular metabolism and morphogenesis.
The degenerate structure of the full enzymatic model suggests rich reduced dynamics near criticality. Exploring rigorous invariant-manifold reductions and more detailed enzyme–substrate formulations is an exciting direction that we are actively pursuing.
8. Acknowledgments
The author warmly thanks Prof. Stepan Timr (J. Heyrovský Institute of Physical Chemistry, Czech Academy of Sciences) for his constant support, stimulating discussions, and many helpful suggestions that greatly improved the quality of this work. FF is also grateful for support from the National Group of Mathematical Physics (GNFM-INdAM).
Statements and Declarations
Competing interests: The author declares that there are no competing interests.
Data availability: No new experimental data were generated in this study. All simulation codes and any numerical data supporting the findings are available from the corresponding author on reasonable request.
Funding: No external funding was received for this research.
Ethics approval and consent to participate: Not applicable.
Authors’ contributions: The author developed the mathematical model, performed all analyses and simulations, and wrote the entire manuscript.